### v1

# Scaling Efficient Masked Autoencoder Learning on Large Remote Sensing Dataset

Abstract
Masked Image Modeling (MIM) has emerged as a pivotal approach for developing foundational visual models in the field of remote sensing (RS). However, current RS datasets are limited in volume and diversity, which significantly constrains the capacity of MIM methods to learn generalizable representations. In this study, we introduce RS-4M, a large-scale dataset designed to enable highly efficient MIM training on RS images. RS-4M comprises 4 million optical images encompassing abundant and fine-grained RS visual tasks, including object-level detection and pixel-level segmentation. Compared to natural images, RS images often contain massive redundant background pixels, which limits the training efficiency of the conventional MIM models. To address this, we propose an efficient MIM method, termed SelectiveMAE, which dynamically encodes and reconstructs a subset of patch tokens selected based on their semantic richness. SelectiveMAE roots in a progressive semantic token selection module, which evolves from reconstructing semantically analogical tokens to encoding complementary semantic dependencies. This approach transforms conventional MIM training into a progressive feature learning process, enabling SelectiveMAE to efficiently learn robust representations of RS images. Extensive experiments show that SelectiveMAE significantly boosts training efficiency by 2.2-2.7 times and enhances the classification, detection, and segmentation performance of the baseline MIM model. The dataset, source code, and trained models will be released at RS-4M.
1 Introduction
Over the past decade, advancements in remote sensing (RS) technology and data acquisition have significantly improved applications in ecosystem monitoring [1], natural disaster management [2], among others [3, 4]. These applications rely on essential capabilities such as scene classification [5, 6], object detection [7], change detection [8], and semantic segmentation [9]. However, each of these downstream tasks often requires substantial computational resources to learn task-specific feature representations and develop specialized models.
Thanks to significant advances in self-supervised learning methods, such as Masked Image Modeling (MIM) techniques [10, 11], the pre-training of visual foundation models has seen remarkable improvements [12, 13, 14, 15, 16, 17]. Consequently, remote sensing foundation models (RSFMs) have recently emerged, offering general feature representations and achieving outstanding performance across various remote sensing downstream tasks [18]. However, two challenges persist in the development of RSFMs. (i) Compared to the ImageNet-21k [19] dataset, previous RS datasets [20, 21, 22, 23] contain significantly fewer samples (approximately 1 million vs. 14 million), which impedes the sufficient MIM training of large backbones. (ii) These datasets, which primarily focus on global scene semantics [20, 21, 22], lack the diversity and fine-grained information of RS scenarios encountered in downstream tasks. This limitation restricts the generalization of the learned representations.
To address these challenges, we introduce a large-scale RS dataset called RS-4M. This dataset, comprising 4 million optical images, is designed to fully leverage the representation learning capabilities of MIM methods in RS applications. RS-4M significantly exceeds previous RS datasets [24, 21, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 18, 37, 38], being at least four times larger. Moreover, RS-4M encompasses a wide range of diverse RS scenarios encountered in downstream tasks such as object-level detection and pixel-level segmentation (Fig. 1 left).
Despite substantial efforts in training RSFM using MIM methods, such as leveraging general image knowledge [25, 26, 28], expanding parameter scales [29], integrating spatio-temporal information [30, 31, 32], and learning multi-scale features [33, 35, 36], the significant computational burden and slow convergence when employing MIM training on large-scale RS datasets cannot be ignored. Specifically, pre-training on 1 million RS samples requires 107 hours for the ViT-B [39] backbone on 8 Nvidia A100 GPUs [18]. In natural scene analysis, this issue has led to numerous studies [40, 41, 42, 43, 44, 45] aimed at improving MIM training efficiency. One approach is to accelerate the token reconstruction process by using decoders with fewer parameters [40, 41]. Another approach is to reduce the number of visible patch tokens input into the vision encoder [42, 43, 44], significantly speeding up feature extraction.
However, conventional MIM approaches, such as the encoding-then-decoding procedure, overlook the unique characteristics of RS images, which typically feature sparse foreground pixels and dense backgrounds [18, 28]. This raises two key questions about how to efficiently conduct MIM training in the RS field: (i) Is it necessary to reconstruct all the redundant background patches during the MIM decoding process? (ii) Is there a feasible way to encode fewer image patches ($\bm{e.g.,\leq}$25%) to accelerate the convergence of MIM training? To address the first question, a measure-based selection process is needed to identify the appropriate patches for reconstruction. For the second question, the intuition is that the patch tokens used in the encoding-then-decoding procedure should effectively capture feature dependencies in RS images.
We propose an efficient MIM method called SelectiveMAE, which dynamically encodes and reconstructs patch tokens based on their semantic richness. Specifically, SelectiveMAE utilizes the Histogram of Oriented Gradients (HOG) algorithm to quantify the semantic richness of patches. Then, it selects a subset of patch tokens ($e.g.,\leq$50%) with higher HOG values for feature encoding ($e.g.,\leq$15%) and pixel reconstruction ($e.g.,\leq$35%). However, using an extremely low ratio of visible patches during MIM training can lead to gradient explosion or vanishing (Fig. 3). To mitigate this, we designed a Progressive Semantic Token Selection (PSTS) module, which dynamically leverages semantically relevant patch tokens throughout the entire MIM training phase. At the beginning, SelectiveMAE encodes semantically rich tokens and reconstructs semantically similar ones to warm up the training process. As training advances, SelectiveMAE shifts to reconstructing high-semantic tokens from encoded lower-semantic ones to capture complementary semantic dependencies. This analogical-to-complementary approach allows SelectiveMAE to efficiently and progressively learn robust representations of RS images, thereby accelerating MIM convergence. Our experiments reveal that only 40% of RS image patches are sufficient to train a comparable MIM model, offering new insights into conducting MIM training on RS images.
The contributions of this paper are summarized as follows. (1) We introduce the RS-4M dataset, a large-scale optical remote sensing dataset for unsupervised learning methods to date, distinguished by its diverse scene details. (2) We propose SelectiveMAE, a novel and efficient MIM method tailored for remote sensing images. This method incorporates a new PSTS module, which significantly accelerates convergence and enhances representation learning compared to the original MIM approach.
Experiment results suggest that compared to the baseline method, SelectiveMAE not only achieves 2.22.7×2.2-2.7$\times$2.2 - 2.7 × speedup in pre-training but also provides performance improvement of 5.6% mAP and 1.4% mF1 when applied to downstream object detection and segmentation tasks, respectively.
2 Related Work
Remote Sensing Datasets. In recent years, many RS datasets have been created for tasks such as scene classification [23, 46], object detection [47, 48, 49], and segmentation [50, 51, 52]. The availability of free, unlabeled satellite images has led to the development of large-scale RS datasets. Some works, like SEN12MS [53] with 180,662 triplets of dual-pol synthetic aperture radar (SAR) and multi-spectral Sentinel-2 image patches, combine various sensor data to create extensive datasets, while many other datasets focus solely on optical RS images. MillionAID [20] offers a million-scale dataset for RS optical image classification, while SeCo [22] and CACo [21] provide nearly a million images of the same location over different times. These datasets primarily target scene classification and often overlook fine-grained target information, limiting their utility for various downstream tasks. To address this gap, we introduce the RS-4M dataset, which is larger and more diverse, enhancing performance across multiple downstream tasks.
Remote Sensing Foundation Models. remains unlabeled and thus inaccessible for supervised learning [54]. Self-supervised learning methods have recently been employed to extract representations from unlabeled RS data. Although there are some contrastive self-supervised methods [24, 21, 22, 55], recent advancements have primarily centered around generative self-supervised methods. For example, RVSA [18] leverages the vision transformer backbone pre-trained by the vanilla MAE [10] method and introduces computationally efficient rotated varied-size window attention to replace the original attention during fine-tuning. In addition, many studies are focusing on improving generative self-supervised algorithms by leveraging general image knowledge [25, 26, 28], scaling up parameter sizes [29], integrating spatio-temporal data [30, 31, 32], handling multi-sensor data [56, 57, 58, 59, 60], and employing multi-scale concepts [33, 35, 36]. However, these methods have not effectively addressed the substantial computational burden associated with self-supervised pre-training in RS. In this paper, we propose a novel acceleration method that significantly speeds up training, enhancing practicality in the RS domain.
Masked Image Modeling. Inspired by the success of Masked Language Modeling (MLM) in NLP [61], MIM has been developed for visual pre-training [62, 11, 10, 63, 64]. MIM learns image representations by reconstructing masked tokens, focusing on various regression targets [65, 66, 67, 68, 69], masking strategies [70, 71], and reconstruction methods [72, 73, 74, 75]. For instance, MAE [10] demonstrates that predicting pixel values can be as effective as using complex targets. A major challenge for MIM is its high computational demand and lengthy pre-training times. To mitigate this, some studies use asymmetric encoder-decoder strategies [40, 41], reduce input patches [42, 44] or ues a novel Difficulty-Flatten Loss [43]. Additionally, CrossMAE [45] employs cross-attention between masked and visible tokens to enhance efficiency without sacrificing performance. However, these methods do not account for the unique characteristics of RS images, such as sparse foreground information and complex backgrounds. In this paper, we introduce a novel PSTS module to address this issue and accelerate the training process.
3 Dataset
3.1 Data collection and organization
Figure 1: Left: The organization of data sources in RS-4M. Right: The comparison between RS-4M and other datasets.
Recent advancements in self-supervised pre-training RSFMs are hindered by the smaller scale and diversity of existing RS datasets compared to natural scene datasets. To address this issue, we propose the RS-4M dataset, a continuously updated large-scale RS dataset following the principles of Diversity, Richness, and Scalability (DiRS) [20]. (1) We collected and organized datasets from the past decade, focusing on mainstream RS image interpretation tasks, including scene classification, object detection, semantic segmentation, and change detection. These datasets, detailed in Fig. 1, encompass around 100 sub-datasets with a total of 13.5 million images. (2) We excluded multispectral and SAR data, keeping only optical images in this study. Future updates will incorporate multispectral and SAR data to support multi-modal self-supervised pre-training. (3) Unlike datasets like MillionAID [20] designed for scene classification, our RS dataset is designed to facilitate downstream detection and segmentation tasks. To address large-scale image in segmentation datasets, we randomly cropped high-resolution images into smaller slices. (4) Images with pixels below 64 or above 1024 were removed to facilitate self-supervised pre-training. (5) We combined the remaining images and eliminated duplicates using a two-phase process: a coarse phase with perceptual hashing[76] and a refined phase with manual review. With the Hash value to select and the checking of manual review, only highly similar images are excluded. This resulted in the RS-4M dataset containing about 4 million high-quality RS optical images, which is four times larger than previous representative RS datasets, as shown in Fig. 1 right.
Table 1: Comparison between RS-4M and MillionAID: ViT Base models [39] are pre-trained on these datasets using MAE [10] and fine-tuned on the RESISC-45 dataset [77].
Dataset Images Number Epoch Top-1 Accuracy
MillionAID [20] 1 Million 800 89.20
RS-4M 2 Million 400 91.80
RS-4M 3 Million 267 92.24
RS-4M 4 Million 200 92.38
3.2 Preliminary Evaluation of RS-4M Dataset
RS-4M offers a significantly larger and more diverse image set compared to previous datasets. To evaluate its effectiveness, we pre-train a ViT base model [39] using the vanilla MAE method [10]. For comparison, we use the MillionAID dataset, maintaining an equal number of data points during training: 800 epochs for MillionAID’s 1 million images and 200 epochs for our RS-4M dataset. In addition, we create new datasets by sampling different numbers of images from RS-4M for further comparison. After pre-training, we fine-tune the models on the downstream RESISC-45 dataset [77] and assess the Top-1 classification accuracy. The results in Table 1 demonstrate that the RS-4M dataset outperforms MillionAID across various sample sizes, with the highest accuracy of 92.38% achieved using the entire dataset. Even with fewer images, RS-4M consistently delivers superior performance, highlighting its greater diversity and effectiveness. These findings underscore the advantage of using larger and more varied datasets for unsupervised pre-training.
4 Method
4.1 Masked Autoencoders Preliminaries
1) Masking. Similar to supervised training of a standard ViT, MAE divides the image into regular, non-overlapping patches. It then samples a subset of these patches and masks the remaining ones. Typically, the masking ratio is 75%, meaning only 25% of the patches are input to the encoder. This random sampling follows a uniform distribution according to the masking ratio. 2) MAE Encoder. The encoder is a standard ViT applied only to the visible, unmasked patches. It linearly projects the patches, adds positional embeddings, and processes them through a series of transformer blocks. By operating on a smaller subset of patches, the encoder enables training of large models with reduced computational and memory requirements. 3) MAE Decoder.The encoded tokens and masked tokens are fed into the decoder, which comprises transformer blocks with self-attention layers. The masked tokens are shared, learnable tensors enhanced with positional embeddings. The decoder, utilized only during pre-training, generates the output predictions for those masked tokens. 4) Reconstruction Target. MAE predicts the pixel values for each masked patch, with each element in the decoder output representing a patch’s pixel value vector. The loss function computes the mean squared error (MSE) between the reconstructed targets and original patches.
RS optical images typically contain many redundant background pixels, which aligns well with MAE’s masking strategy. In MAE, the encoder processes only 25% of the patches, significantly reducing computational load by avoiding processing the entire image. Building on MAE, we aim to leverage the redundancy in RS images to accelerate training. Specifically, we address two questions: 1) Is it necessary to reconstruct all the masked patches given the redundancy in RS images? 2) Can the visible patches input to the MAE encoder be further compressed to enhance acceleration?
Figure 2: Overview of SelectiveMAE. The vanilla version of SelectiveMAE.
4.2 Partial Reconstruction
Table 2: Results of different partial reconstruction strategies. Reconstruction Ratio=25%. Top-1 classification accuracy on RESISC-45 [77] is reported.
Method Selection Data Throu. Acc.
/Min.
Partial Reco. False 17k 89.2
Partial Reco. Random 25k 88.8
Partial Reco. ATS [78] 20k 85.7
Partial Reco. HOG 25k 89.2
For question 1, previous research [45] has shown that for general images, when MAE reconstructs 75% of the patches to calculate the loss, a specially designed decoder doesn’t need to fully reconstruct all remaining patches. In fact, reconstructing just 50% or even 25% of the patches can achieve similar performance and speed up training. However, for RS images, if we randomly sample patches and remove most for reconstruction, the reconstructed patches might not be semantically rich ones. As shown in Table 2, using only a random subset for reconstruction degrades performance.
To address this issue, we propose selecting semantically rich patches for reconstruction instead of random selection. Specifically, given an input image $x\in\mathbb{R}^{H\times W\times C}$, it is reshaped into $N=(H\times W)/p^{2}$ non-overlapping patches $x^{p}\in\mathbb{R}^{N\times(p^{2}C)}$, where $p$ is the patch size, $(H,W)$ is the size of the input image, and $C$ is the number of channels. These patches $\{x_{i}^{p}\}_{i=1}^{N}$ are then linearly mapped to patch embeddings. To retain positional information, positional embeddings are added to the patches. We select a portion of the patches to input to the encoder based on the masking ratio $m\in[0,1]$ ($m=85\%$ by default), as detailed in Sec. 4.1. The remaining patches serve as reconstruction targets for the decoder. Unlike MAE’s masking ratio, we introduce a new reconstruction ratio $r$, the proportion of pixels to be reconstructed, denoted as $r\in[0,m]$ ($r=25\%$ by default). We compute the HOG features $HOG(\cdot)$ of the remaining patches and select those with the high HOG feature values according to the reconstruction ratio $r$, rather than using all patches. The process can be formulated as:
[…]
where $token_{R}$ denotes the selected mask tokens for reconstruction and $\text{top}_{n}(\cdot)$ denotes the index set of the selected top $n$ tokens. The decoder uses a lightweight design based on cross-attention following CrossMAE [45]. Experimental results in Table 2 show that this partial reconstruction strategy significantly increases the training throughput without affecting the learned representations.
4.3 Progressive Semantic Token Selection
For the second question, we initially tried a naive approach by increasing the masking ratio to 85%, meaning only 15% of the patches in each RS image were input to the encoder while keeping a 25% reconstruction ratio as proposed in Sec. 4.2. However, during training, this setup often led to issues like gradient explosions or loss divergence, as shown in Fig. 3. To illustrate, we continued MAE’s training despite these gradient explosions. The figure shows that using only 40% (=15%+25%) of the patches for encoding and reconstruction frequently caused unstable training due to gradient explosions. We speculate that, although RS images have background redundancy, using such a small portion of patches significantly challenges MAE training.
Figure 3: Left: Baseline training loss. Using only 40% of masked patches for reconstruction often leads to gradient explosions. Right: Training loss using PSTS.
How can we achieve acceleration at a high masking ratio (e.g., 85%) while ensuring MAE completes the pre-training task for RS images? Some studies suggest that machine learning algorithms often benefit from starting with simpler tasks and gradually increasing the difficulty, a strategy known as curriculum learning [79, 80, 81]. This approach involves progressively introducing more challenging examples during training. While curriculum learning typically focuses on whole image samples, self-supervised training, which masks large portions of image pixels, emphasizes patch-level selection. Therefore, in self-supervised training, it’s essential to distinguish between “easy" and “hard" patches.
To this end, we introduce the Progressive Semantic Token Selection (PSTS) module for patch selection. In this module, depicted in Fig. 2, we begin by selecting a limited number of patches and then select additional patches based on them in the training epoch, dynamically transitioning from easily learned, semantically similar patches to more challenging, complementary ones.
For initialization, we employ a HOG selection strategy to choose the initial patch set from $S^{N}=\left\{x_{i}^{p}\right\}_{i=1}^{N}$, with a proportion $s\in[0,(1-m)/2]$. We define the initial set of selected token as:
[…]
We select $\lfloor\ s\times N\rfloor$ tokens with the maximum HOG feature values (i.e., ${top}_{\lfloor\ s\times N\rfloor}$) from the original token set to form the initial token set. This simple yet effective strategy ensures that semantically rich tokens are selected.
Following token initialization, we aim to incrementally increase the number of tokens to facilitate a training progression from easy to difficult, while adhering to the final masking ratio for selected tokens. This process is outlined in Algorithm 1. Given the initial token sets’ high HOG feature values, nearby tokens selected by PSTS also exhibit high HOG feature values. Recalling our partial reconstruction approach in Section 4.2, we filter the remaining unselected token set to include those with high HOG values as reconstruction targets. Consequently, when selecting nearby tokens, the tokens input for encoding and those chosen for reconstruction are closer, thereby simplifying the training process. Conversely, if a token’s feature distribution significantly deviates from others in the set, indicating it is farther away, we consider such patches more challenging to learn.
Specifically, we select tokens from $S^{U}$ based on $S^{I}$. First, we use $\mathbb{S^{I}}\in\mathbb{R}^{|S^{K}|\times d}$ and $\mathbb{S^{U}}\in\mathbb{R}^{|S^{U}|\times d}$ to denote the matrix representation of the initial token set $S^{I}$ and the unselected token set $S^{U}$, where ||$\cdot$|| ⋅ | represents the number of tokens and $d$ the feature dimension after the embedding layer. We use Cosine Distance to measure the distance between the tokens in these two sets:
$\mathcal{D}(S^{U},S^{I})=\mathbf{1}-cos\left\langle S^{U},S^{I}\right\rangle=\mathbf{1}-\frac{\mathbb{S^{U}}\mathbb{S^{I}}^{\mathrm{T}}}{\left|\kern-1.07639pt\left|\mathbb{S^{U}}\right|\kern-1.07639pt\right|\cdot\left|\kern-1.07639pt\left|\mathbb{S^{I}}\right|\kern-1.07639pt\right|},$ (3)
where $\mathbf{1}$ is an all-one matrix. $\mathcal{D}(S^{U},S^{I})\in\mathbb{R}^{|S^{U}|\times|S^{I}|}$ represents the pairwise distances between tokens in $S^{U}$ and $S^{I}$. Next, we define the distance between the tokens in $S^{U}$ to the initial token set $S^{K}$ based on the selection criteria in each training stage as follows:
[…]
where $i\in\{1,\cdots,|S^{U}|\}$, $j\in\{1,\cdots,|S^{I}|\}$, and $\zeta$ represents the training stage depending on the number of epochs. Finally, we sample $\lfloor N\times(1-m-s)\rfloor$ tokens from ${S^{U}}$ and add them with ${S^{I}}$ to form ${S^{K}}$, which can be formulated as follows:
[…]
${S^{K}}={S^{I}}\cup\mit{S}^{*},\quad{S^{U}}={S^{U}}\setminus\mit{S}^{*},$ (6)
where $\mit{S}^{*}$ represents the selected token from ${S^{U}}$. The operations in Eq. (5) and Eq. (6) are performed in each training epoch. This process is summarized in Algorithm 1.
Algorithm 1 Progressive Semantic Token Selection
1:Number of training epochs $T$ and total training stages $N_{g}$ (i.e., $T/N_{g}$ epochs for each stage), masking rate $m$, input dataset $\mathcal{X}$
2:Obtain the selected tokens set ${S^{K}}$ and update ${S^{U}}$ in each epoch
3:for $t\leftarrow 1$ to $T$ do
4:     Sample data sample from $\mathcal{X}$, feed-forwarded through the embedding to obtain the output token set ${S}_{N}$
5:     $s\leftarrow\frac{1}{2}(1-m)$
6:     Obtain ${S}^{I}$ via Eq. (2) and initialize ${S^{U}}$
7:     Obtain the current training stage $\zeta=\lceil N_{g}*t/T\rceil$
[…]
9:     Obtain $S^{*}$ and update $S^{K},S^{U}$ via Eq. (5) and Eq. (6)
10:end for
5 Experiment
In this section, we first discuss the pre-training process using the RS-4M dataset, comparing the training speed and performance of our proposed SelectiveMAE with existing MAE acceleration methods. We then demonstrate SelectiveMAE’s superior performance across various downstream tasks compared to different RSFMs. Finally, we explore design choices such as the reconstruction ratio, mask ratio, and decoder depth through extensive ablation experiments, detailed in the Appendix.
5.1 Pre-training
Pre-training Setup. Our pre-training experiment setup largely follows MAE [10]. Unlike other MIM methods in the RS field, such as SatMAE [30] and ScaleMAE [33], we train both the commonly used ViT-B and the larger ViT-L backbones [39]. We adjust the input image to a size of $224\times 224$, with a patch size of 16, using the AdamW optimizer and a cosine learning rate scheduler. The initial learning rate is set to 1.5e-4, batch size to 1024, masking ratio to 85%, and reconstruction ratio to 25%. For patch selection during the 800 epochs, the adjustment thresholds are set at 300 epochs, with the last 200 epochs using random distance to enhance robustness.
Table 3: Top-1 classification accuracy on RESISC-45 [77]. “-” indicates unavailable accuracy data. Training throughput was measured using 8 NVIDIA 40G A100 GPUs.
Model Publication Backbone Params Data Throughput/Min. Acccuracy
MAE [10] CVPR’22 ViT-B 86M 17k 93.45
CrossMAE [45] Arxiv’24 ViT-B 86M 22k 92.82
SatMAE [30] NIPS’22 ViT-B 86M 15k -
ScaleMAE [33] ICCV’23 ViT-B 86M 16k -
SelectiveMAE - ViT-B 86M 37k ($\times2$) 94.58 ($+1.13$)
MAE [10] CVPR’22 ViT-L 307M 15k 93.80
SatMAE [30] NIPS’22 ViT-L 307M 13k 94.10
ScaleMAE [33] ICCV’23 ViT-L 307M 13k 95.04
SelectiveMAE - ViT-L 307M 41k (2.7×2.7$\times$2.7 ×) 95.77 ($+1.97$)
Pre-training Results.
We pre-trained ViT-B and ViT-L on the RS-4M dataset, evaluating the image data throughput per minute for SelectiveMAE and the baseline (MAE) [10]. We also assessed the training speed of the general domain acceleration method CrossMAE [45], as well as the remote sensing methods SatMAE [30] and ScaleMAE [33]. Table 3 illustrates our method’s more than doubled acceleration for both base and large ViT models compared to the vanilla MAE baseline. It also outperforms the general domain acceleration method CrossMAE significantly. Additionally, we reported the top-1 classification accuracy on the same dataset, with both MAE and CrossMAE trained on MillionAID [20]. Table 3 clearly demonstrates our method’s superior speed and performance.
The significant acceleration effect of SelectiveMAE stems from utilizing only 40% of the image patches (15% for the encoder and 25% for the decoder). Despite this reduced utilization, it achieves superior performance compared to other acceleration methods [45] due to the proposed novel token masking and reconstruction strategy, tailored for the remote sensing images. This unique approach ensures that SelectiveMAE delivers both speed and high performance, as validated in Table 4.
5.2 Fine-tuning on Downstream Tasks
We further evaluated SelectiveMAE across three key downstream tasks: scene classification, object detection, and semantic segmentation. In addition to benchmarking against MAE-based approaches like SatMAE[30], ScaleMAE[33], SSL4EO[82], RVSA [18], RingMo[37], and GFM[25], we compared it with other leading visual foundation models. These encompass methods leveraging contrastive learning and enhanced backbones, such as GASSL [24], SeCo [22], TOV [28], and CACo [21], and advanced supervised learning based approaches [83]. Notably, none of these comparative methods were specifically tailored for accelerated pre-training, and although their speed performance wasn’t reported, we anticipate they may lag behind the computationally efficient baseline (MAE). To ensure fairness, all experiments followed to consistent fine-tuning settings. Supplementary materials and code provide comprehensive experimental details.
Table 4: Results for scene classification, object detection, and semantic segmentation. “TR" represents the ratio of training data to the entire dataset.
Model Publication Backbone Sence Object Semantic
Classification Detection Segmentation
AID RESISC-45 DIOR DIOR-R LoveDA SpaceNetv1
(TR=20%/50%) (TR=10%/20%) mAP50 mAP50 mIoU mF1
SeCo [22] ICCV’21 ResNet-50 93.47/95.99 89.64/92.91 - - 43.63 77.09
GASSL [24] ICCV’21 ResNet-50 93.55/95.92 90.86/93.06 67.40 65.65 48.76 78.51
TOV [28] JSTARS’23 ResNet-50 95.16/97.09 90.97/93.79 70.16 66.33 49.70 -
CACo [21] CVPR’23 ResNet-50 90.88/95.05 88.28/91.94 66.91 64.10 48.89 77.94
SatMAE [30] NIPS’22 ViT-L 95.02/96.94 91.72/94.10 70.89 65.66 - 78.07
ScaleMAE [33] ICCV’23 ViT-L 96.44/97.58 92.63/95.04 73.81 66.47 - -
SSL4EO [82] GRSM’23 ViT-S 91.06/94.74 87.60/91.27 64.82 61.23 - -
RingMo [37] TGRS’22 Swin-B 96.90/98.34 94.25/95.67 75.90 - - -
SatLas [83] ICCV’23 Swin-B 94.96/97.38 92.16/94.70 74.10 67.59 - -
GFM [25] ICCV’23 Swin-B 95.47/97.09 92.73/94.64 72.84 67.67 - -
RVSA [18] TGRS’23 ViT-B+RVSA 97.03/98.50 93.93/95.69 75.80 68.06 51.95 -
SelectiveMAE - ViT-B 96.78/98.12 93.35/94.58 75.70 67.78 53.05 79.50
SelectiveMAE - ViT-L 97.25/98.48 94.57/95.77 77.80 70.31 54.31 79.46
Scene Classification. We begin by evaluating the pre-trained model’s performance on the scene classification task, which offers insight into its overall representation capability without requiring additional decoders. We leverage two scene classification datasets: AID [84] and RESISC-45 [77]. Training details, including the training and test split ratio, adhere to [18, 33], with further specifics available in our code. Evaluation is based on overall accuracy (OA). Table 4 presents the results, demonstrating SelectiveMAE’s competitive performance compared to other pre-training methods on both datasets. Notably, our method enjoys a significantly faster training speed, e.g., over 2$\times$ faster than the MAE baseline.
Besides, it outperforms RVSA and others when scaled to ViT-L. This demonstrates that, with our SelectiveMAE training method on the RS-4M dataset, the model can learn strong feature representations and scale efficiently.
Horizontal & Oriented Object Detection. We utilized the well-established DIOR dataset for horizontal object detection [85] and its improved variant DIOR-R for oriented object detection [86], both comprising RGB images. Following the methodologies of previous works [18, 37], we maintained consistent experimental setups, employing Faster-RCNN [87] and Oriented-RCNN [88] as detectors for each dataset. Detailed experimental configurations are available in the accompanying code. The results are summarized in Table 4. Our approach, utilizing a ViT-B backbone, demonstrates competitive or superior performance compared to other methods with a Swin-B backbone, such as RingMo [37], SatLas [83], and GFM [25]. This is particularly noteworthy as hierarchical transformers generally excel at generating multiscale features essential for object detection. Moreover, our SelectiveMAE with a ViT-B backbone outperforms SatMAE [30] and ScaleMAE [33], which use a larger ViT-L backbone, and matches the advanced RVSA [18, 34] that employs specialized window attention. When using the larger ViT-L backbone, SelectiveMAE shows significantly enhanced performance across both detection datasets, underscoring the excellent scalability of our method.
Semantic Segmentation. We further evaluate the performance of the pre-trained model on pixel-level perception tasks, particularly semantic segmentation, using two well-known remote sensing datasets: LoveDA [51] and SpaceNetv1 [52]. Our implementation closely follows [18], utilizing UperNet [89] as the segmentation framework, with mean Intersection over Union (mIOU) as the evaluation metric. Table 4 demonstrates the clear superiority of SelectiveMAE over its competitors in pixel-level semantic segmentation. The selection and reconstruction strategies employed by SelectiveMAE focus on semantically rich patches, such as the boundaries between foreground objects and background stuff, leading to better representation learning for segmentation.
6 Conclusion
In this paper, we introduce RS-4M, a large-scale optical remote sensing dataset for unsupervised learning. Unlike previous datasets in the RS field, RS-4M offers a significantly larger and more diverse image set with fine-grained details relevant to downstream tasks. Benchmarking representative MIM methods on RS-4M highlights its advantages in these tasks. To address the computational burden of MIM training on large-scale RS datasets, we propose SelectiveMAE, an efficient MIM method that dynamically encodes and reconstructs patch tokens based on their semantic richness. SelectiveMAE significantly accelerates training, demonstrating that only 40% of RS image patches are needed to train a comparable MIM model. Experiments show that SelectiveMAE not only achieves over 2×2$\times$2 × speedup in pre-training but also enhances performance in various downstream tasks.
Limitations and Discussion. Although the RS-4M dataset surpasses existing RS datasets by more than four times, its size can be further increased by collecting and filtering additional RS data. Additionally, its diversity can be enhanced by incorporating data from various types of sensors.
    A: The dataset was created by:
    - Fengxiang Wang (National University of Defense Technology),
    - Hongzhen Wang (Tsinghua University),
    - Di Wang (Wuhan University),
    - Zonghao Guo (University of Chinese Academic of Sciences),
    - Zhenyu Zhong (Nankai University),
    - Long Lan (National University of Defense Technology),
    - Jing Zhang (The University of Sydney),
    - Zhiyuan Liu (Tsinghua University),
    - Maosong Sun (Tsinghua University).
  3. 3.
    “Who funded the creation of the dataset?”
    A: The dataset creation was funded by the affiliations of the authors involved in this work.
7.5.2 Composition
Most of the questions in this section are intended to provide dataset consumers with the information they need to make informed decisions about using the dataset for their chosen tasks. Some of the questions are designed to elicit information about compliance with the EU’s General Data Protection Regulation (GDPR) or comparable regulations in other jurisdictions. Questions that apply only to datasets that relate to people are grouped together at the end of the section. We recommend taking a broad interpretation of whether a dataset relates to people. For example, any dataset containing text that was written by people relates to people.
  1. 1.
    “What do the instances that comprise our datasets represent (e.g., documents, photos, people, countries)?”
    A: The dataset primarily comprises visible light remote sensing images captured by satellites. All datasets utilized in RS-4M are publicly accessible and nonprofit.
  2. 2.
    “How many instances are there in total (of each type, if appropriate)?”
    A: RS-4M contains 4 million remote sensing image instances captured by satellites.
  3. 3.
    “Does the dataset contain all possible instances or is it a sample (not necessarily random) of instances from a larger set?”
    A: Yes, our dataset contains all possible instances that have been collected so far.
  4. 4.
    “Is there a label or target associated with each instance?”
    A: No, our dataset is intended for self-supervised learning. Therefore, each instance is an individual remote sensing image and does not contain annotations.
  5. 5.
    “Is any information missing from individual instances?”
    A: No.
  6. 6.
    “Are relationships between individual instances made explicit (e.g., users’ movie ratings, social network links)?”
    A: Yes, the relationship between individual instances is explicit.
  7. 7.
    “Are there recommended data splits (e.g., training, development/validation, testing)?”
    A: Yes, the entire dataset is intended for self-supervised methods, and we recommend using the whole dataset for self-supervised learning research.
  8. 8.
    “Is the dataset self-contained, or does it link to or otherwise rely on external resources (e.g., websites, tweets, other datasets)?”
    A: Yes, our dataset relies on many publicly available remote sensing datasets, which we have detailed in the main text.
  9. 9.
    “Does the dataset contain data that might be considered confidential (e.g., data that is protected by legal privilege or by doctor–patient confidentiality, data that includes the content of individuals’ non-public communications)?”
    A: No, all data are clearly licensed.
  10. 10.
    “Does the dataset contain data that, if viewed directly, might be offensive, insulting, threatening, or might otherwise cause anxiety?”
    A: No.
7.5.3 Collection Process
In addition to the goals outlined in the previous section, the questions in this section are designed to elicit information that may help researchers and practitioners create alternative datasets with similar characteristics. Again, questions that apply only to datasets that relate to people are grouped together at the end of the section.
  1. 1.
    “How was the data associated with each instance acquired?”
    A: Please refer to the details listed in the main text Sec. 3.
  2. 2.
    “What mechanisms or procedures were used to collect the data (e.g., hardware apparatuses or sensors, manual human curation, software programs, software APIs)?”
    A: Please refer to the details listed in the main text Sec. 3.
  3. 3.
    “If the dataset is a sample from a larger set, what was the sampling strategy (e.g., deterministic, probabilistic with specific sampling probabilities)?”
    A: Please refer to the details listed in the main text Sec. 3.
7.5.4 Preprocessing, Cleaning, and Labeling
The questions in this section are intended to provide dataset consumers with the information they need to determine whether the “raw” data has been processed in ways that are compatible with their chosen tasks. For example, text that has been converted into a “bag-of-words" is not suitable for tasks involving word order.
  1. 1.
    “Was any preprocessing/cleaning/labeling of the data done (e.g., discretization or bucketing, tokenization, part-of-speech tagging, SIFT feature extraction, removal of instances, processing of missing values)?”
    A: Yes, we preprocessed and cleaned data in our dataset.
  2. 2.
    “Was the ‘raw’ data saved in addition to the preprocessed/cleaned/labeled data (e.g., to support unanticipated future uses)?”
    A: Yes, raw data is accessible.
  3. 3.
    “Is the software that was used to preprocess/clean/label the data available?”
    A: Yes, the necessary software used to preprocess and clean the data is publicly available.
7.5.5 Uses
The questions in this section are intended to encourage dataset creators to reflect on tasks for which the dataset should and should not be used. By explicitly highlighting these tasks, dataset creators can help dataset consumers make informed decisions, thereby avoiding potential risks or harms.
  1. 1.
    “Has the dataset been used for any tasks already?”
    A: No.
  2. 2.
    “Is there a repository that links to any or all papers or systems that use the dataset?”
    A: Yes, we provide such links in our GitHub repository.
  3. 3.
    “What (other) tasks could the dataset be used for?”
    A: The dataset could be used for training the remote sensing foundation models with the self-supervised learning method.
  4. 4.
    “Is there anything about the composition of the dataset or the way it was collected and preprocessed/cleaned/labeled that might impact future uses?”
    A: N/A.
  5. 5.
    “Are there tasks for which the dataset should not be used?”
    A: N/A.
7.5.6 Distribution
Dataset creators should provide answers to these questions prior to distributing the dataset either internally within the entity on behalf of which the dataset was created or externally to third parties.
  1. 1.
    “Will the dataset be distributed to third parties outside of the entity (e.g., company, institution, organization) on behalf of which the dataset was created?”
    A: No.
  2. 2.
    “How will the dataset be distributed (e.g., tarball on website, API, GitHub)?”
    A: Very likely to be distributed by website, API, and GitHub repository.
  3. 3.
    “When will the dataset be distributed?”
    A: The datasets are publicly accessible.
  4. 4.
    “Will the dataset be distributed under a copyright or other intellectual property (IP) license, and/or under applicable terms of use (ToU)?”
    A: Yes, the dataset is under the Creative Commons Attribution-NonCommercial-ShareAlike 4.0 International License.
  5. 5.
    “Have any third parties imposed IP-based or other restrictions on the data associated with the instances?”
    A: No.
  6. 6.
    “Do any export controls or other regulatory restrictions apply to the dataset or to individual instances?”
    A: No.
7.5.7 Maintenance
As with the questions in the previous section, dataset creators should provide answers to these questions prior to distributing the dataset. The questions in this section are intended to encourage dataset creators to plan for dataset maintenance and communicate this plan to dataset consumers.
  1. 1.
    “Who will be supporting/hosting/maintaining the dataset?”
    A: The authors of this work serve to support, host, and maintain the datasets.
  2. 2.
    “How can the owner/curator/manager of the dataset be contacted (e.g., email address)?”
    A: The curators can be contacted via the email addresses listed on our webpage [^4] https://github.com/Fengxiang23/SelectiveMAE..
  3. 3.
    “Is there an erratum?”
    A: There is no explicit erratum; updates and known errors will be specified in future versions.
  4. 4.
    “Will the dataset be updated (e.g., to correct labeling errors, add new instances, delete instances)?”
    A: Yes, for the current version. Future updates (if any) will be posted on the dataset website.
  5. 5.
    “Will older versions of the dataset continue to be supported/hosted/maintained?”
    A: Yes. This is the first version of the release; future updates will be posted and older versions will be replaced.
  6. 6.
    “If others want to extend/augment/build on/contribute to the dataset, is there a mechanism for them to do so?”
    A: Yes, we provide detailed instructions for future extensions.
7.6 Limitation and Potential Societal Impact
In this section, we elaborate on the limitations and potential societal impact of this work.
7.6.1 Potential Limitations
While RS-4M provides a comprehensive benchmark for training the remote sensing foundation models with self-supervised learning methods, there are several limitations to consider:
  • Scope of Sensors: Although our benchmark includes 4 million visible light remote sensing images, it may not cover all possible real-world scenarios. There could be additional sensor data, like multispectral data that were not included in this study, potentially limiting the generalizability of our findings.
  • Model and Dataset Diversity: While our dataset is primarily focused on adapting to downstream tasks like detection and segmentation, there is undeniably less data available for these tasks compared to general scene classification. In the future, we should collect more data that is better suited for various downstream tasks.
  • Computation and Resource Requirements: Pre-training on the extensive RS-4M dataset, comprising 4 million images, demands substantial computational resources, despite our introduction of the efficient SelectiveMAE method. This may limit access to the benchmark for research groups lacking ample computational power.
7.6.2 Potential Negative Societal Impact
While the development of remote sensing foundation models with self-supervised learning methods has the potential to significantly advance remote sensing downstream tasks, there are potential negative societal impacts that must be considered:
  • Safety Risks: Our benchmark aims to improve remote sensing foundation models, but relying too heavily on these models may breed overconfidence in autonomous systems. It’s crucial to deploy these systems with adequate safety measures and human oversight to uphold public safety.
  • Environmental Impact: Training and evaluating remote sensing foundation models demand substantial computational resources, leading to a notable environmental footprint. Encouraging the adoption of energy-efficient algorithms (i.e., the proposed SelectiveMAE) and sustainable computing practices is crucial to reduce the environmental impact of this research.
  • Bias and Fairness: The performance of foundational remote sensing models can vary across different environments and conditions, potentially introducing biases in downstream tasks within the remote sensing domain. It is essential to train and evaluate these models on diverse datasets to ensure fairness and avoid discriminatory outcomes.

### v3

@@
-Scaling Efficient Masked Autoencoder Learning on Large Remote Sensing Dataset
+OpticalRS-4M: Scaling Efficient Masked Autoencoder Learning on Large Remote Sensing Dataset
 Abstract
-Masked Image Modeling (MIM) has emerged as a pivotal approach for developing foundational visual models in the field of remote sensing (RS).
-However, current RS datasets are limited in volume and diversity, which significantly constrains the capacity of MIM methods to learn generalizable representations.
-In this study, we introduce RS-4M, a large-scale dataset designed to enable highly efficient MIM training on RS images.
-RS-4M comprises 4 million optical images encompassing abundant and fine-grained RS visual tasks, including object-level detection and pixel-level segmentation.
-Compared to natural images, RS images often contain massive redundant background pixels, which limits the training efficiency of the conventional MIM models.
-To address this, we propose an efficient MIM method, termed SelectiveMAE, which dynamically encodes and reconstructs a subset of patch tokens selected based on their semantic richness.
-SelectiveMAE roots in a progressive semantic token selection module, which evolves from reconstructing semantically analogical tokens to encoding complementary semantic dependencies.
-This approach transforms conventional MIM training into a progressive feature learning process, enabling SelectiveMAE to efficiently learn robust representations of RS images.
-Extensive experiments show that SelectiveMAE significantly boosts training efficiency by 2.2-2.7 times and enhances the classification, detection, and segmentation performance of the baseline MIM model.
-The dataset, source code, and trained models will be released at RS-4M.
+Masked Image Modeling (MIM) has become an essential method for building foundational visual models in remote sensing (RS).
+However, the limitations in size and diversity of existing RS datasets restrict the ability of MIM methods to learn generalizable representations.
+Additionally, conventional MIM techniques, which require reconstructing all tokens, introduce unnecessary computational overhead.
+To address these issues, we present a new pre-training pipeline for RS models, featuring the creation of a large-scale RS dataset and an efficient MIM approach.
+We curated a high-quality dataset named OpticalRS-4M by collecting publicly available RS datasets and processing them through exclusion, slicing, and deduplication.
+OpticalRS-4M comprises 4 million optical images covering various RS tasks, such as object detection and pixel segmentation.
+To enhance efficiency, we propose SelectiveMAE, a pre-training method that dynamically encodes and reconstructs semantically rich patch tokens, thereby reducing the inefficiencies of traditional MIM models caused by redundant background pixels in RS images.
+Extensive experiments demonstrate that OpticalRS-4M significantly improves classification, detection, and segmentation performance, while SelectiveMAE increases training efficiency over 2$\times$ times.
+This highlights the effectiveness and scalability of our pipeline in developing RS foundational models.
+The dataset, source code, and trained models will be released at OpticalRS-4M.
 1 Introduction
-Over the past decade, advancements in remote sensing (RS) technology and data acquisition have significantly improved applications in ecosystem monitoring [1], natural disaster management [2], among others [3, 4].
+Over the past decade, advancements in remote sensing (RS) technology and data acquisition have improved applications in ecosystem monitoring [1], natural disaster management [2], among others [3, 4].
 These applications rely on essential capabilities such as scene classification [5, 6], object detection [7], change detection [8], and semantic segmentation [9].
 However, each of these downstream tasks often requires substantial computational resources to learn task-specific feature representations and develop specialized models.
-Thanks to significant advances in self-supervised learning methods, such as Masked Image Modeling (MIM) techniques [10, 11], the pre-training of visual foundation models has seen remarkable improvements [12, 13, 14, 15, 16, 17].
+Thanks to advances in self-supervised learning methods, such as Masked Image Modeling (MIM) techniques [10, 11], the pre-training of visual foundation models has seen remarkable improvements [12, 13, 14, 15, 16, 17].
 Consequently, remote sensing foundation models (RSFMs) have recently emerged, offering general feature representations and achieving outstanding performance across various remote sensing downstream tasks [18].
 However, two challenges persist in the development of RSFMs.
-(i) Compared to the ImageNet-21k [19] dataset, previous RS datasets [20, 21, 22, 23] contain significantly fewer samples (approximately 1 million vs.
+(i) Compared to the ImageNet-21k [19] dataset, previous RS datasets [20, 21, 22, 23] contain fewer samples (approximately 1 million vs.
 14 million), which impedes the sufficient MIM training of large backbones.
@@
 This limitation restricts the generalization of the learned representations.
-To address these challenges, we introduce a large-scale RS dataset called RS-4M.
-This dataset, comprising 4 million optical images, is designed to fully leverage the representation learning capabilities of MIM methods in RS applications.
-RS-4M significantly exceeds previous RS datasets [24, 21, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 18, 37, 38], being at least four times larger.
-Moreover, RS-4M encompasses a wide range of diverse RS scenarios encountered in downstream tasks such as object-level detection and pixel-level segmentation (Fig.
+To address these challenges, we propose a new pipeline to collect, create, and efficiently process a large optical RS dataset.
+Firstly, we reviewed publicly available remote sensing datasets from the past decade and selected them based on the DiRS (Diversity, Richness, and Scalability) principle [20].
+Nevertheless, issues such as inconsistent data sources, excessively large image sizes, and redundant pixels still exist.
+Therefore, we also apply exclusion, slicing, and deduplication processes to further improve data quality.
+Correspondingly, we obtain a large-scale RS dataset called OpticalRS-4M comprising 4 million optical images, which is designed to fully leverage the representation learning capabilities of MIM methods in RS applications.
+OpticalRS-4M exceeds previous RS datasets [24, 21, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 18, 37, 38], being at least four times larger.
+Moreover, OpticalRS-4M encompasses a wide range of diverse RS scenarios encountered in downstream tasks such as object-level detection and pixel-level segmentation (Fig.
 1 left).
-Despite substantial efforts in training RSFM using MIM methods, such as leveraging general image knowledge [25, 26, 28], expanding parameter scales [29], integrating spatio-temporal information [30, 31, 32], and learning multi-scale features [33, 35, 36], the significant computational burden and slow convergence when employing MIM training on large-scale RS datasets cannot be ignored.
+Despite substantial efforts in training RSFM using MIM methods, such as leveraging general image knowledge [25, 26, 28], expanding parameter scales [29], integrating spatio-temporal information [30, 31, 32], and learning multi-scale features [33, 35, 36], the computational burden and slow convergence when employing MIM training on large-scale RS datasets cannot be ignored.
 Specifically, pre-training on 1 million RS samples requires 107 hours for the ViT-B [39] backbone on 8 Nvidia A100 GPUs [18].
+This issue becomes even more pronounced when training on larger datasets, e.g., OpticalRS-4M.
 In natural scene analysis, this issue has led to numerous studies [40, 41, 42, 43, 44, 45] aimed at improving MIM training efficiency.
 One approach is to accelerate the token reconstruction process by using decoders with fewer parameters [40, 41].
-Another approach is to reduce the number of visible patch tokens input into the vision encoder [42, 43, 44], significantly speeding up feature extraction.
+Another approach is to reduce the number of visible patch tokens input into the vision encoder [42, 43, 44], speeding up feature extraction.
 However, conventional MIM approaches, such as the encoding-then-decoding procedure, overlook the unique characteristics of RS images, which typically feature sparse foreground pixels and dense backgrounds [18, 28].
-This raises two key questions about how to efficiently conduct MIM training in the RS field: (i) Is it necessary to reconstruct all the redundant background patches during the MIM decoding process?
-(ii) Is there a feasible way to encode fewer image patches ($\bm{e.g.,\leq}$25%) to accelerate the convergence of MIM training?
+This raises two key questions about how to efficiently conduct MIM training in the RS field: 1) Is it necessary to reconstruct all the redundant background patches during the MIM decoding process?
+2) Is there a feasible way to encode fewer image patches ($\bm{e.g.,\leq}$25%) to accelerate the convergence of MIM training?
 To address the first question, a measure-based selection process is needed to identify the appropriate patches for reconstruction.
 For the second question, the intuition is that the patch tokens used in the encoding-then-decoding procedure should effectively capture feature dependencies in RS images.
-We propose an efficient MIM method called SelectiveMAE, which dynamically encodes and reconstructs patch tokens based on their semantic richness.
+Regarding the above issues, in the second part of the pipeline, we introduce a MIM method called SelectiveMAE for efficiently processing RS images, which dynamically encodes and reconstructs patch tokens based on their semantic richness.
 Specifically, SelectiveMAE utilizes the Histogram of Oriented Gradients (HOG) algorithm to quantify the semantic richness of patches.
 Then, it selects a subset of patch tokens ($e.g.,\leq$50%) with higher HOG values for feature encoding ($e.g.,\leq$15%) and pixel reconstruction ($e.g.,\leq$35%).
-However, using an extremely low ratio of visible patches during MIM training can lead to gradient explosion or vanishing (Fig.
-3).
-To mitigate this, we designed a Progressive Semantic Token Selection (PSTS) module, which dynamically leverages semantically relevant patch tokens throughout the entire MIM training phase.
-At the beginning, SelectiveMAE encodes semantically rich tokens and reconstructs semantically similar ones to warm up the training process.
+However, using an extremely low ratio of visible patches during MIM training can lead to gradient explosion or vanishing.
+To mitigate this, we designed a Progressive Semantic Token Selection (PSTS) module, which dynamically selects semantically relevant patch tokens during the entire training phase.
+In the beginning, SelectiveMAE encodes semantically rich tokens and reconstructs semantically similar ones to warm up the training process.
 As training advances, SelectiveMAE shifts to reconstructing high-semantic tokens from encoded lower-semantic ones to capture complementary semantic dependencies.
-This analogical-to-complementary approach allows SelectiveMAE to efficiently and progressively learn robust representations of RS images, thereby accelerating MIM convergence.
-Our experiments reveal that only 40% of RS image patches are sufficient to train a comparable MIM model, offering new insights into conducting MIM training on RS images.
-The contributions of this paper are summarized as follows.
-(1) We introduce the RS-4M dataset, a large-scale optical remote sensing dataset for unsupervised learning methods to date, distinguished by its diverse scene details.
-(2) We propose SelectiveMAE, a novel and efficient MIM method tailored for remote sensing images.
-This method incorporates a new PSTS module, which significantly accelerates convergence and enhances representation learning compared to the original MIM approach.
-Experiment results suggest that compared to the baseline method, SelectiveMAE not only achieves 2.22.7×2.2-2.7$\times$2.2 - 2.7 × speedup in pre-training but also provides performance improvement of 5.6% mAP and 1.4% mF1 when applied to downstream object detection and segmentation tasks, respectively.
+This analogical-to-complementary strategy allows SelectiveMAE to efficiently and progressively learn robust representations of RS images while accelerating MIM convergence.
+Our experiments reveal that 40% of RS image patches are sufficient to train a comparable model, offering new insights into MIM training on RS images.
+The contributions of this paper are summarized as follows:
+1) We introduce a new pipeline to collect, create, and efficiently process a large optical RS dataset for developing RS foundation models.
+2) Using this pipeline, we create the OpticalRS-4M dataset, which is a large-scale RS dataset comprising 4 million optical images with diversified coverage scenarios.
+3) We introduce SelectiveMAE, an efficient MIM method tailored for RS image pre-training.
+It significantly accelerates convergence and enhances representation learning compared to the original MIM approach.
+Experiment results suggest the effectiveness and scalability of the proposed pipeline.
+OpticalRS-4M significantly enhances the performance of RS foundation models in downstream tasks, while SelectiveMAE achieves over 2×2$\times$2 × speedup in pre-training compared to MAE [10].
 2 Related Work
@@
 The availability of free, unlabeled satellite images has led to the development of large-scale RS datasets.
-Some works, like SEN12MS [53] with 180,662 triplets of dual-pol synthetic aperture radar (SAR) and multi-spectral Sentinel-2 image patches, combine various sensor data to create extensive datasets, while many other datasets focus solely on optical RS images.
+Some works combine various sensor data to create extensive datasets.
+SEN12MS [53] with 180,662 triplets of dual-pol synthetic aperture radar (SAR) and multi-spectral Sentinel-2 image patches.
+SSL4EO-S12 [54] dataset uses publicly available SAR/optical satellite data collected by the European Space Agency’s Sentinel mission.
+SSL4EO-L [55] consists entirely of multispectral data.
+SatlasPretrain [56] is constructed from data sources such as Sentinel-2, Sentinel-1, Landsat 8/9, and 0.5 - 2 meter/pixel aerial imagery.
+While it includes millions of images, the majority are multispectral and SAR, with only a small fraction being RGB.
+There are also some large-scale RS visual-language datasets such as Skyscript [57] and RS5M [58], but they primarily focus on multimodal tasks.
+Currently, some datasets focus solely on optical RS images.
 MillionAID [20] offers a million-scale dataset for RS optical image classification, while SeCo [22] and CACo [21] provide nearly a million images of the same location over different times.
-These datasets primarily target scene classification and often overlook fine-grained target information, limiting their utility for various downstream tasks.
-To address this gap, we introduce the RS-4M dataset, which is larger and more diverse, enhancing performance across multiple downstream tasks.
+However, these datasets primarily target scene classification and often overlook fine-grained target information, limiting their utility for various downstream tasks.
+To address this gap, with the proposed pipeline, we introduce the OpticalRS-4M dataset, which is larger and more diverse, enhancing performance across multiple downstream tasks.
+Masked Image Modeling.
+Inspired by the success of Masked Language Modeling (MLM) in NLP [59], MIM has been developed for visual pre-training [60, 11, 10, 61, 62].
+MIM learns image representations by reconstructing masked tokens, focusing on various regression targets [63, 64, 65, 66, 67], masking strategies [68, 69], and reconstruction methods [70, 71, 72, 73].
+For instance, MAE [10] demonstrates that predicting pixel values can be as effective as using complex targets.
+A major challenge for MIM is its high computational demand and lengthy pre-training times.
+To mitigate this, some studies use asymmetric encoder-decoder strategies [40, 41], reduce input patches [42, 44] or use a novel Difficulty-Flatten Loss [43].
+Additionally, CrossMAE [45] employs cross-attention between masked and visible tokens to enhance efficiency without sacrificing performance.
+However, these methods do not account for the unique characteristics of RS images, such as sparse foreground information and complex backgrounds.
+Considering these issues, we introduce an efficient MIM method named SelectiveMAE that significantly speeds up pre-training, enhancing the practicality of developing RS foundation models based on large-scale datasets.
 Remote Sensing Foundation Models.
-remains unlabeled and thus inaccessible for supervised learning [54].
+Despite the abundance of RS data, much of it remains unlabeled and thus inaccessible for supervised learning [74].
 Self-supervised learning methods have recently been employed to extract representations from unlabeled RS data.
@@
 However, these methods have not effectively addressed the substantial computational burden associated with self-supervised pre-training in RS.
-In this paper, we propose a novel acceleration method that significantly speeds up training, enhancing practicality in the RS domain.
-Masked Image Modeling.
-Inspired by the success of Masked Language Modeling (MLM) in NLP [61], MIM has been developed for visual pre-training [62, 11, 10, 63, 64].
-MIM learns image representations by reconstructing masked tokens, focusing on various regression targets [65, 66, 67, 68, 69], masking strategies [70, 71], and reconstruction methods [72, 73, 74, 75].
-For instance, MAE [10] demonstrates that predicting pixel values can be as effective as using complex targets.
-A major challenge for MIM is its high computational demand and lengthy pre-training times.
-To mitigate this, some studies use asymmetric encoder-decoder strategies [40, 41], reduce input patches [42, 44] or ues a novel Difficulty-Flatten Loss [43].
-Additionally, CrossMAE [45] employs cross-attention between masked and visible tokens to enhance efficiency without sacrificing performance.
-However, these methods do not account for the unique characteristics of RS images, such as sparse foreground information and complex backgrounds.
-In this paper, we introduce a novel PSTS module to address this issue and accelerate the training process.
-3 Dataset
-3.1 Data collection and organization
-Figure 1: Left: The organization of data sources in RS-4M. Right: The comparison between RS-4M and other datasets.
-Recent advancements in self-supervised pre-training RSFMs are hindered by the smaller scale and diversity of existing RS datasets compared to natural scene datasets.
-To address this issue, we propose the RS-4M dataset, a continuously updated large-scale RS dataset following the principles of Diversity, Richness, and Scalability (DiRS) [20].
-(1) We collected and organized datasets from the past decade, focusing on mainstream RS image interpretation tasks, including scene classification, object detection, semantic segmentation, and change detection.
-These datasets, detailed in Fig.
-1, encompass around 100 sub-datasets with a total of 13.5 million images.
-(2) We excluded multispectral and SAR data, keeping only optical images in this study.
-Future updates will incorporate multispectral and SAR data to support multi-modal self-supervised pre-training.
-(3) Unlike datasets like MillionAID [20] designed for scene classification, our RS dataset is designed to facilitate downstream detection and segmentation tasks.
-To address large-scale image in segmentation datasets, we randomly cropped high-resolution images into smaller slices.
-(4) Images with pixels below 64 or above 1024 were removed to facilitate self-supervised pre-training.
-(5) We combined the remaining images and eliminated duplicates using a two-phase process: a coarse phase with perceptual hashing[76] and a refined phase with manual review.
-With the Hash value to select and the checking of manual review, only highly similar images are excluded.
-This resulted in the RS-4M dataset containing about 4 million high-quality RS optical images, which is four times larger than previous representative RS datasets, as shown in Fig.
+In the paper, we propose a new pipeline that can collect, create, and efficiently process large amounts of optical RS data for developing RS foundation models, enhancing the practicality of MIM pre-training on large-scale datasets.
+3 Methodology
+The proposed pipeline contains two critical components: dataset generation and efficient pre-training, which will be introduced in detail in the following text.
+3.1 Dataset Generation
+Figure 1: Left: The organization of data sources in OpticalRS-4M. Right: The comparison between OpticalRS-4M and other datasets.
+Recent progress in self-supervised pre-training for RSFMs is limited by the relatively small scale and diversity of existing RS datasets compared to natural scene datasets.
+To overcome this, we first curate a large-scale RS dataset, OpticalRS-4M, through extensive data collection and preprocessing.
+3.1.1 Data Collection
+We began by reviewing publicly available RS datasets from the past decade.
+Recognizing that not all data are suitable for self-supervised pre-training, we adopted specific criteria for data collection, as outlined in DiRS [20]: 1) Diversity: A dataset is diverse if its images capture various typical visual features of relevant scenes or targets and offer complementarity.
+Intra-class Diversity: High diversity within classes ensures comprehensive representation of real-world target distributions.
+To ensure this, we consider the data from different sources of the same scene or target.
+Moreover, in the RS field, the attributes of objects may vary with geographic location and imaging time, so the change detection data is also incorporated.
+Inter-class Diversity: To enable self-supervised pre-training to distinguish between categories effectively, we incorporated as many fine-grained categories as possible, especially those with high semantic overlap.
+This strategy helps the model learn robust feature representations.
+2) Richness: Beyond diversity, richness is also essential.
+We ensure diverse content features and large sample sizes by collecting images under various conditions — weather, seasons, lighting, and so on — introducing variability in translation, viewpoint, object posture, spatial resolution, background, and occlusion.
+Our data collection process actively includes the sample of the same category under different imaging conditions to ensure comprehensive real-world representation, enhancing the model’s representation and generalization capabilities.
+3) Scalability: Given the evolving applications of remote sensing images, scalability is also crucial.
+We constantly enrich the dataset based on scene and target categories across classification, detection, and segmentation tasks.
+In summary, following the guidelines, we constructed a large-scale RS image dataset with diverse coverage scenarios. Fig. 1 left illustrates the included RS datasets collected.
+3.1.2 Data Preprocessing
+Following data collection, challenges such as inconsistent data sources, oversized images, and redundant pixels remained.
+To address these, we implemented a standardized, scalable data preprocessing workflow consisting of four steps:
+1) We focused exclusively on optical images for this study, excluding multispectral and SAR data, though future updates will incorporate these modalities for multi-modal self-supervised pre-training.
+2) To manage the large image sizes, we randomly cropped high-resolution images into smaller slices.
+3) We filtered out images with fewer than 64 or more than 1024 pixels.
+4) The remaining images were combined and duplicates were removed using a two-phase approach: first, a coarse phase with perceptual hashing [81], and then a refined phase involving manual review.
+This ensured that only highly similar images were excluded.
+As a result of these steps, we created the OpticalRS-4M dataset, which contains approximately 4 million high-quality RS images — significantly larger than previous representative RS datasets, as illustrated in Fig.
 1 right.
-Table 1: Comparison between RS-4M and MillionAID: ViT Base models [39] are pre-trained on these datasets using MAE [10] and fine-tuned on the RESISC-45 dataset [77].
-Dataset Images Number Epoch Top-1 Accuracy
-MillionAID [20] 1 Million 800 89.20
-RS-4M 2 Million 400 91.80
-RS-4M 3 Million 267 92.24
-RS-4M 4 Million 200 92.38
-3.2 Preliminary Evaluation of RS-4M Dataset
-RS-4M offers a significantly larger and more diverse image set compared to previous datasets.
-To evaluate its effectiveness, we pre-train a ViT base model [39] using the vanilla MAE method [10].
-For comparison, we use the MillionAID dataset, maintaining an equal number of data points during training: 800 epochs for MillionAID’s 1 million images and 200 epochs for our RS-4M dataset.
-In addition, we create new datasets by sampling different numbers of images from RS-4M for further comparison.
-After pre-training, we fine-tune the models on the downstream RESISC-45 dataset [77] and assess the Top-1 classification accuracy.
-The results in Table 1 demonstrate that the RS-4M dataset outperforms MillionAID across various sample sizes, with the highest accuracy of 92.38% achieved using the entire dataset.
-Even with fewer images, RS-4M consistently delivers superior performance, highlighting its greater diversity and effectiveness.
-These findings underscore the advantage of using larger and more varied datasets for unsupervised pre-training.
-4 Method
-4.1 Masked Autoencoders Preliminaries
+In summary, we followed the DiRS principles for data collection, applying exclusion, slicing, and deduplication to ensure high-quality results.
+We’ve screened and processed most small and medium-scale datasets from the past decade in this field.
+More efforts can be made to extend OpticalRS-4M by including larger datasets like SSL4EO-S12 [54] and SSL4EO-L [55] to enhance dataset capacity.
+3.1.3 Dataset Diversity
+Refer to captionRefer to caption
+Figure 2: Left: Main categories with visual examples from OpticalRS-4M.Right: Full hierarchy of OpticalRS-4M categories.
+To demonstrate the broad scope of scenes and target categories covered by our dataset, we conducted a statistical analysis accompanied by a detailed label explanation, reviewing all the scenarios and targets included in OpticalRS-4M, as shown in Fig.
+2.
+The analysis reveals that OpticalRS-4M comprises 12 main categories, each containing numerous subclasses.
+Unlike existing self-supervised optical remote sensing datasets, OpticalRS-4M not only offers more comprehensive scene and target information but also introduces “Events” as a distinct high-level category.
+This category includes labels such as “Concert”, “Parade/Protest”, “Religious Activity”, “Fire”, “Flood”, “Landslide”, “Post-Earthquake”, and “Traffic Collision”.
+This expanded information enhances the versatility of OpticalRS-4M for a wider range of downstream tasks.
+3.2 Efficient Pre-training
+After obtaining a large-scale RS dataset, the next step is to conduct the self-supervised pre-training.
+However, as the capacity of the dataset grows, vast computational and time costs are required for existing MIM approaches used in the RS community, e.g., MAE [10].
+To address this issue, our pipeline introduces an efficient MIM method, i.e., SelectiveMAE.
+In this section, we first review the preliminaries of MAE, then we present the details of SelectiveMAE.
+3.2.1 Masked Autoencoders Preliminaries
 1) Masking.
@@
 It linearly projects the patches, adds positional embeddings, and processes them through a series of transformer blocks.
-By operating on a smaller subset of patches, the encoder enables training of large models with reduced computational and memory requirements.
+By operating on a smaller subset of patches, the encoder enables the training of large models with reduced computational and memory requirements.
 3) MAE Decoder.The encoded tokens and masked tokens are fed into the decoder, which comprises transformer blocks with self-attention layers.
@@
 RS optical images typically contain many redundant background pixels, which aligns well with MAE’s masking strategy.
-In MAE, the encoder processes only 25% of the patches, significantly reducing computational load by avoiding processing the entire image.
+In MAE, the encoder processes only 25% of the patches, reducing computational load by avoiding processing the entire image.
 Building on MAE, we aim to leverage the redundancy in RS images to accelerate training.
@@
 Figure 3: Overview of SelectiveMAE. The vanilla version of SelectiveMAE.
-4.2 Partial Reconstruction
-Table 2: Results of different partial reconstruction strategies. Reconstruction Ratio=25%. Top-1 classification accuracy on RESISC-45 [77] is reported.
-Method Selection Data Throu. Acc.
-/Min.
-Partial Reco. False 17k 89.2
-Partial Reco. Random 25k 88.8
-Partial Reco. ATS [78] 20k 85.7
-Partial Reco. HOG 25k 89.2
+3.2.2 Partial Reconstruction
 For question 1, previous research [45] has shown that for general images, when MAE reconstructs 75% of the patches to calculate the loss, a specially designed decoder doesn’t need to fully reconstruct all remaining patches.
@@
 However, for RS images, if we randomly sample patches and remove most for reconstruction, the reconstructed patches might not be semantically rich ones.
-As shown in Table 2, using only a random subset for reconstruction degrades performance.
+Using only a random subset for reconstruction even degrades performance.
 To address this issue, we propose selecting semantically rich patches for reconstruction instead of random selection.
@@
 We select a portion of the patches to input to the encoder based on the masking ratio $m\in[0,1]$ ($m=85\%$ by default), as detailed in Sec.
-4.1.
+3.2.1.
 The remaining patches serve as reconstruction targets for the decoder.
@@
 The decoder uses a lightweight design based on cross-attention following CrossMAE [45].
-Experimental results in Table 2 show that this partial reconstruction strategy significantly increases the training throughput without affecting the learned representations.
-4.3 Progressive Semantic Token Selection
+Experimental results in the Appendix show that this partial reconstruction strategy significantly increases the training throughput without affecting the learned representations.
+3.2.3 Progressive Semantic Token Selection
 For the second question, we initially tried a naive approach by increasing the masking ratio to 85%, meaning only 15% of the patches in each RS image were input to the encoder while keeping a 25% reconstruction ratio as proposed in Sec.
-4.2.
-However, during training, this setup often led to issues like gradient explosions or loss divergence, as shown in Fig.
-3.
-To illustrate, we continued MAE’s training despite these gradient explosions.
-The figure shows that using only 40% (=15%+25%) of the patches for encoding and reconstruction frequently caused unstable training due to gradient explosions.
-We speculate that, although RS images have background redundancy, using such a small portion of patches significantly challenges MAE training.
-Figure 3: Left: Baseline training loss. Using only 40% of masked patches for reconstruction often leads to gradient explosions. Right: Training loss using PSTS.
-How can we achieve acceleration at a high masking ratio (e.g., 85%) while ensuring MAE completes the pre-training task for RS images?
-Some studies suggest that machine learning algorithms often benefit from starting with simpler tasks and gradually increasing the difficulty, a strategy known as curriculum learning [79, 80, 81].
-This approach involves progressively introducing more challenging examples during training.
-While curriculum learning typically focuses on whole image samples, self-supervised training, which masks large portions of image pixels, emphasizes patch-level selection.
-Therefore, in self-supervised training, it’s essential to distinguish between “easy" and “hard" patches.
+3.2.2.
+However, during training, this setup often led to issues like gradient explosions or loss divergence (see the Appendix).
+We speculate that, although RS images have background redundancy, using such a small portion of patches challenges MAE training.
+How can we achieve acceleration at a high masking ratio while ensuring MAE convergence?
+Curriculum learning [82, 83, 84] may offer a solution.
+However, it typically focuses on selecting whole image samples, whereas MIM pre-training masks large portions of image pixels, emphasizing patch-level selection.
+Therefore, it’s essential to distinguish between “easy” and “hard” patches in MIM pre-training.
 To this end, we introduce the Progressive Semantic Token Selection (PSTS) module for patch selection.
 In this module, depicted in Fig.
-2, we begin by selecting a limited number of patches and then select additional patches based on them in the training epoch, dynamically transitioning from easily learned, semantically similar patches to more challenging, complementary ones.
-For initialization, we employ a HOG selection strategy to choose the initial patch set from $S^{N}=\left\{x_{i}^{p}\right\}_{i=1}^{N}$, with a proportion $s\in[0,(1-m)/2]$.
-We define the initial set of selected token as:
-[…]
-We select $\lfloor\ s\times N\rfloor$ tokens with the maximum HOG feature values (i.e., ${top}_{\lfloor\ s\times N\rfloor}$) from the original token set to form the initial token set.
+3, we begin by selecting a limited number of patches and then select additional patches based on them in the training epoch, dynamically transitioning from easily learned, semantically similar patches to more challenging, complementary ones.
+For initialization, we employ a HOG selection strategy to choose the initial patch set.
 This simple yet effective strategy ensures that semantically rich tokens are selected.
 Following token initialization, we aim to incrementally increase the number of tokens to facilitate a training progression from easy to difficult, while adhering to the final masking ratio for selected tokens.
-This process is outlined in Algorithm 1.
+This process is outlined in the Appendix.
 Given the initial token sets’ high HOG feature values, nearby tokens selected by PSTS also exhibit high HOG feature values.
-Recalling our partial reconstruction approach in Section 4.2, we filter the remaining unselected token set to include those with high HOG values as reconstruction targets.
+Recalling our partial reconstruction approach in Sec.
+3.2.2, we filter the remaining unselected token set to include those with high HOG values as reconstruction targets.
 Consequently, when selecting nearby tokens, the tokens input for encoding and those chosen for reconstruction are closer, thereby simplifying the training process.
-Conversely, if a token’s feature distribution significantly deviates from others in the set, indicating it is farther away, we consider such patches more challenging to learn.
-Specifically, we select tokens from $S^{U}$ based on $S^{I}$.
-First, we use $\mathbb{S^{I}}\in\mathbb{R}^{|S^{K}|\times d}$ and $\mathbb{S^{U}}\in\mathbb{R}^{|S^{U}|\times d}$ to denote the matrix representation of the initial token set $S^{I}$ and the unselected token set $S^{U}$, where ||$\cdot$|| ⋅ | represents the number of tokens and $d$ the feature dimension after the embedding layer.
-We use Cosine Distance to measure the distance between the tokens in these two sets:
-$\mathcal{D}(S^{U},S^{I})=\mathbf{1}-cos\left\langle S^{U},S^{I}\right\rangle=\mathbf{1}-\frac{\mathbb{S^{U}}\mathbb{S^{I}}^{\mathrm{T}}}{\left|\kern-1.07639pt\left|\mathbb{S^{U}}\right|\kern-1.07639pt\right|\cdot\left|\kern-1.07639pt\left|\mathbb{S^{I}}\right|\kern-1.07639pt\right|},$ (3)
-where $\mathbf{1}$ is an all-one matrix.
-$\mathcal{D}(S^{U},S^{I})\in\mathbb{R}^{|S^{U}|\times|S^{I}|}$ represents the pairwise distances between tokens in $S^{U}$ and $S^{I}$.
-Next, we define the distance between the tokens in $S^{U}$ to the initial token set $S^{K}$ based on the selection criteria in each training stage as follows:
-[…]
-where $i\in\{1,\cdots,|S^{U}|\}$, $j\in\{1,\cdots,|S^{I}|\}$, and $\zeta$ represents the training stage depending on the number of epochs.
-Finally, we sample $\lfloor N\times(1-m-s)\rfloor$ tokens from ${S^{U}}$ and add them with ${S^{I}}$ to form ${S^{K}}$, which can be formulated as follows:
-[…]
-${S^{K}}={S^{I}}\cup\mit{S}^{*},\quad{S^{U}}={S^{U}}\setminus\mit{S}^{*},$ (6)
-where $\mit{S}^{*}$ represents the selected token from ${S^{U}}$. The operations in Eq. (5) and Eq. (6) are performed in each training epoch. This process is summarized in Algorithm 1.
-Algorithm 1 Progressive Semantic Token Selection
-1:Number of training epochs $T$ and total training stages $N_{g}$ (i.e., $T/N_{g}$ epochs for each stage), masking rate $m$, input dataset $\mathcal{X}$
-2:Obtain the selected tokens set ${S^{K}}$ and update ${S^{U}}$ in each epoch
-3:for $t\leftarrow 1$ to $T$ do
-4:     Sample data sample from $\mathcal{X}$, feed-forwarded through the embedding to obtain the output token set ${S}_{N}$
-5:     $s\leftarrow\frac{1}{2}(1-m)$
-6:     Obtain ${S}^{I}$ via Eq. (2) and initialize ${S^{U}}$
-7:     Obtain the current training stage $\zeta=\lceil N_{g}*t/T\rceil$
-[…]
-9:     Obtain $S^{*}$ and update $S^{K},S^{U}$ via Eq. (5) and Eq. (6)
-10:end for
+Conversely, if a token’s feature distribution deviates from others in the set, indicating it is farther away, we consider such patches more challenging to learn.
+Details of the process are presented in the Appendix.
 4 Experiment
-In this section, we first discuss the pre-training process using the RS-4M dataset, comparing the training speed and performance of our proposed SelectiveMAE with existing MAE acceleration methods.
-We then demonstrate SelectiveMAE’s superior performance across various downstream tasks compared to different RSFMs.
-Finally, we explore design choices such as the reconstruction ratio, mask ratio, and decoder depth through extensive ablation experiments, detailed in the Appendix.
-5.1 Pre-training
+In this section, we first discuss the effectiveness of the dataset and acceleration strategy in the pipeline through ablation experiments.
+We also conducted further experiments to test the scalability of the pipeline in the ablation experiments.
+Additionally, we performed comprehensive comparative experiments to evaluate the pipeline’s performance on downstream tasks.
 Pre-training Setup.
 Our pre-training experiment setup largely follows MAE [10].
-Unlike other MIM methods in the RS field, such as SatMAE [30] and ScaleMAE [33], we train both the commonly used ViT-B and the larger ViT-L backbones [39].
-We adjust the input image to a size of $224\times 224$, with a patch size of 16, using the AdamW optimizer and a cosine learning rate scheduler.
-The initial learning rate is set to 1.5e-4, batch size to 1024, masking ratio to 85%, and reconstruction ratio to 25%.
-For patch selection during the 800 epochs, the adjustment thresholds are set at 300 epochs, with the last 200 epochs using random distance to enhance robustness.
-Table 3: Top-1 classification accuracy on RESISC-45 [77]. “-” indicates unavailable accuracy data. Training throughput was measured using 8 NVIDIA 40G A100 GPUs.
-Model Publication Backbone Params Data Throughput/Min. Acccuracy
-MAE [10] CVPR’22 ViT-B 86M 17k 93.45
-CrossMAE [45] Arxiv’24 ViT-B 86M 22k 92.82
-SatMAE [30] NIPS’22 ViT-B 86M 15k -
-ScaleMAE [33] ICCV’23 ViT-B 86M 16k -
-SelectiveMAE - ViT-B 86M 37k ($\times2$) 94.58 ($+1.13$)
-MAE [10] CVPR’22 ViT-L 307M 15k 93.80
-SatMAE [30] NIPS’22 ViT-L 307M 13k 94.10
-ScaleMAE [33] ICCV’23 ViT-L 307M 13k 95.04
-SelectiveMAE - ViT-L 307M 41k (2.7×2.7$\times$2.7 ×) 95.77 ($+1.97$)
-Pre-training Results.
-We pre-trained ViT-B and ViT-L on the RS-4M dataset, evaluating the image data throughput per minute for SelectiveMAE and the baseline (MAE) [10].
-We also assessed the training speed of the general domain acceleration method CrossMAE [45], as well as the remote sensing methods SatMAE [30] and ScaleMAE [33].
-Table 3 illustrates our method’s more than doubled acceleration for both base and large ViT models compared to the vanilla MAE baseline.
-It also outperforms the general domain acceleration method CrossMAE significantly.
-Additionally, we reported the top-1 classification accuracy on the same dataset, with both MAE and CrossMAE trained on MillionAID [20].
-Table 3 clearly demonstrates our method’s superior speed and performance.
-The significant acceleration effect of SelectiveMAE stems from utilizing only 40% of the image patches (15% for the encoder and 25% for the decoder).
-Despite this reduced utilization, it achieves superior performance compared to other acceleration methods [45] due to the proposed novel token masking and reconstruction strategy, tailored for the remote sensing images.
-This unique approach ensures that SelectiveMAE delivers both speed and high performance, as validated in Table 4.
-5.2 Fine-tuning on Downstream Tasks
-We further evaluated SelectiveMAE across three key downstream tasks: scene classification, object detection, and semantic segmentation.
+The masking ratio is 85%, and the reconstruction ratio is 25%.
+For patch selection, the adjustment threshold is set at 300 epochs, with the last 200 epochs using random distance to enhance robustness.
+4.1 Ablation Study
+To validate the pipeline’s effectiveness, we conducted separate evaluations of both the OpticalRS-4M dataset generated by the pipeline and the accelerated training method, SelectiveMAE.
+Additionally, we assessed the pipeline’s scalability in collecting and processing datasets of varying sizes.
+4.1.1 Superiority of OpticalRS-4M
+Table 1: Comparison of downstream task performance at matched data throughput between MillionAID and OpticalRS-4M, assessed using MAE [10] pre-training and ViT Base models [39].
+Dataset Method Images Number Epoch Sence Object Semantic
+Classification Detection Segmentation
+AID RESISC-45 DIOR DIOR-R LoveDA SpaceNetv1
+OA (TR=20% / 50%) OA (TR=10% / 20%) $\mathrm{mAP}_{50}$ $\mathrm{mAP}_{50}$ mIoU mF1
+MillionAID MAE 1 million 800 94.92/97.38 89.20/93.60 71.80 62.33 51.24 79.24
+OpticalRS-4M MAE 1 million 800 96.26/97.98 91.53/93.88 75.90 66.46 53.10 79.35
+OpticalRS-4M MAE 2 million 400 96.64/98.10 91.80/94.31 73.90 65.95 52.86 79.37
+OpticalRS-4M MAE 3 million 267 96.67/98.18 92.24/94.41 75.40 67.07 52.39 79.37
+OpticalRS-4M MAE 4 million 200 96.10/98.03 92.38/94.30 74.70 66.26 52.75 79.23
+Dataset Diversity.
+In self-supervised pre-training of RSFMs, the primary goal is to enhance downstream task performance.
+We define dataset diversity as its ability to improve various types of these tasks.
+While datasets like MillionAID focus on classification, our OpticalRS-4M dataset is designed to extend the benefit to detection and segmentation tasks.
+We assess this by evaluating the effectiveness of models trained on different datasets with equivalent data throughput.
+For instance, MillionAID, with 1 million samples, is trained using the basic MAE method for 800 epochs as a benchmark.
+To showcase the diversity of OpticalRS-4M, containing 4 million samples, we train it with MAE for just 200 epochs.
+Additionally, we validate OpticalRS-4M’s diversity by progressively reducing its training data — down to 3 million, 2 million, or even 1 million samples — while keeping data throughput constant.
+As shown in Table 1, our experiments demonstrate that even when randomly sampling 1 million images from OpticalRS-4M, it outperforms MillionAID in downstream tasks, particularly in object detection and segmentation.
+This consistent superior performance, even with fewer images, highlights OpticalRS-4M’s greater diversity and effectiveness, emphasizing the benefits of using larger and more varied datasets for unsupervised pre-training.
+The results in Table 1 indicate that the optimal outcomes at equivalent data throughput are achieved with 3 million samples rather than 4 million.
+In most cases, the model pre-trained on 3 million samples outperforms across various downstream tasks.
+Notably, on the DIOR and LoveDA datasets, the best results were obtained with 1 million samples trained for 800 epochs.
+However, as shown in Table 2, when we extended the training on 4 million samples to more epochs, the model achieved the highest average performance across most downstream tasks.
+This suggests that training on 4 million samples for only 200 epochs was insufficient for the MIM method to fully leverage the large dataset, resulting in inferior generalization in downstream tasks.
+Table 2: Performance on downstream tasks with extended training on OpticalRS-4M.
+Dataset Method Images Number Epoch Sence Object Semantic
+Classification Detection Segmentation
+AID RESISC-45 DIOR DIOR-R LoveDA SpaceNetv1
+OA (TR=20% / 50%) OA (TR=10% / 20%) $\mathrm{mAP}_{50}$ $\mathrm{mAP}_{50}$ mIoU mF1
+OpticalRS-4M MAE 3 million 267 96.67/98.18 92.24/94.41 75.40 67.07 52.39 79.37
+OpticalRS-4M MAE 4 million 100 95.12/97.54 91.26/93.85 - - - -
+OpticalRS-4M MAE 4 million 200 96.10/98.03 92.38/94.30 74.70 66.26 52.75 79.23
+OpticalRS-4M MAE 4 million 400 96.36/98.16 92.41/94.03 - - - -
+OpticalRS-4M MAE 4 million 800 96.88/98.22 92.44/94.43 75.40 67.35 52.80 79.41
+Extended training on OpticalRS-4M.
+To maximize the potential of OpticalRS-4M, we extended training using the MAE pre-training method and ViT-B, varying only the number of training epochs while keeping other settings consistent.
+The fine-tuning results, shown in Table 2, reveal that as the number of epochs increases, OpticalRS-4M shows greater advantages in downstream tasks.
+This indicates that extended training is necessary to fully harness OpticalRS-4M’s capabilities.
+However, the gains diminish with additional epochs, suggesting that merely increasing epochs may not suffice.
+To unlock the full potential of OpticalRS-4M, exploring more effective pre-training methods, such as the proposed SelectiveMAE, may be crucial.
+Table 3: Comparison of downstream task performance using SelectiveMAE and MAE [10]. The training time is measured using 8 NVIDIA A100 GPUs.
+Method Dataset Epoch Total Training Sence Object Semantic
+Classification Detection Segmentation
+Hours AID RESISC-45 DIOR DIOR-R LoveDA SpaceNetv1
+OA (TR=20% / 50%) OA (TR=10% / 20%) $\mathrm{mAP}_{50}$ $\mathrm{mAP}_{50}$ mIoU mF1
+MAE OpticalRS-4M 800 177 96.88/98.22 92.44/94.43 75.40 67.35 52.80 79.41
+SelectiveMAE OpticalRS-4M 800 86 (2.1×2.1$\times$2.1 ×) 96.78/98.12 93.35/94.58 75.70 67.78 53.05 79.50
+4.1.2 Effectiveness of SelectiveMAE
+Performance Advantage.
+To further validate the accelerated pre-training strategy of our pipeline, we conducted extensive experiments on OpticalRS-4M using both MAE [10] and SelectiveMAE.
+The performance of ViT-B on various downstream tasks after pre-training is presented in Table 3.
+The results indicate that after pre-training on OpticalRS-4M, SelectiveMAE not only outperforms MAE but also achieves this at 2.1 times the speed.
+This makes SelectiveMAE particularly well-suited for large-scale datasets, offering significant time savings during training on OpticalRS-4M.
+Table 4: The accelerating performance of SelectiveMAE for different backbones compared to MAE [10]. The training time is measured using 8 NVIDIA A100 GPUs.
+Model Backbone Params Total Training Hours
+MAE ViT-B 86M 101
+SelectiveMAE ViT-B 86M 48 (2.1×2.1$\times$2.1 ×)
+MAE ViT-L 307M 113
+SelectiveMAE ViT-L 307M 50 (2.3×2.3$\times$2.3 ×)
+Pre-training Efficiency.
+To further highlight the efficiency advantage of SelectiveMAE, we evaluate the pretraining time with different backbones on the MillionAID dataset, as shown in Table 4.
+It can be seen that, since only 40% of the image patches (15% for the encoder and 25% for the decoder) are involved in the calculation, our methods present more than doubled acceleration for both base and large ViT models compared to the vanilla MAE baseline, where the acceleration advantage is more significant for larger sized models.
+4.1.3 Scalability of the Pipeline
+Table 5: Training cost (measured using 8 NVIDIA A100 GPUs) of SelectiveMAE and MAE [10] on OpticalRS-4M with different data volumes.
+Method Images Number Epoch Total Training Hours
+MAE 1 million 800 54
+SelectiveMAE 1 million 800 25 (-29)
+MAE 4 million 800 177
+SelectiveMAE 4 million 800 86 (-81)
+To evaluate the scalability of our pipeline, we organized datasets containing 1 million and 4 million samples and conducted experiments on each.
+The results, presented in Table 5, show that with 1 million samples, the proposed SelectiveMAE method reduced training time by 29 hours.
+When the dataset was expanded to 4 million samples, the time savings increased to 81 hours.
+These findings demonstrate that our pipeline scales effectively with larger datasets, offering significant pre-training acceleration.
+4.2 Main Results
+4.2.1 Fine-tuning Performance on Downstream Tasks
+We further evaluated SelectiveMAE with the OpticalRS-4M dataset across three key downstream tasks: scene classification, object detection, and semantic segmentation.
 In addition to benchmarking against MAE-based approaches like SatMAE[30], ScaleMAE[33], SSL4EO[54], RVSA [18], RingMo[37], and GFM[25], we compared it with other leading visual foundation models.
-These encompass methods leveraging contrastive learning and enhanced backbones, such as GASSL [24], SeCo [22], TOV [28], and CACo [21], and advanced supervised learning based approaches [83].
+These encompass methods leveraging contrastive learning and enhanced backbones, such as GASSL [24], SeCo [22], TOV [28], and CACo [21], and advanced supervised learning-based approaches [56].
 Notably, none of these comparative methods were specifically tailored for accelerated pre-training, and although their speed performance wasn’t reported, we anticipate they may lag behind the computationally efficient baseline (MAE).
-To ensure fairness, all experiments followed to consistent fine-tuning settings.
-Supplementary materials and code provide comprehensive experimental details.
+More details can be found in the Appendix.
+Scene Classification.
+We begin by evaluating the pre-trained model’s performance on the scene classification task, which offers insight into its overall representation capability without requiring additional decoders.
+We leverage two scene classification datasets: AID [85] and RESISC-45 [86].
+Training details, including the training and test split ratio, adhere to [18, 33], with further specifics available in our code.
+Evaluation is based on overall accuracy (OA).
+Table 6 presents the results, demonstrating SelectiveMAE’s competitive performance compared to other pre-training methods on both datasets.
+Notably, our method enjoys a significantly faster training speed, e.g., over 2$\times$ faster than the MAE baseline.
+Besides, it outperforms RVSA and others when scaled to ViT-L.
+This demonstrates that, with our SelectiveMAE training method on the OpticalRS-4M dataset, the model can learn strong feature representations and scale efficiently.
 Table 6: Results for scene classification, object detection, and semantic segmentation. “TR" represents the ratio of training data to the entire dataset.
@@
 SelectiveMAE - ViT-L 97.25/98.48 94.57/95.77 77.80 70.31 54.31 79.46
-Scene Classification.
-We begin by evaluating the pre-trained model’s performance on the scene classification task, which offers insight into its overall representation capability without requiring additional decoders.
-We leverage two scene classification datasets: AID [84] and RESISC-45 [77].
-Training details, including the training and test split ratio, adhere to [18, 33], with further specifics available in our code.
-Evaluation is based on overall accuracy (OA).
-Table 4 presents the results, demonstrating SelectiveMAE’s competitive performance compared to other pre-training methods on both datasets.
-Notably, our method enjoys a significantly faster training speed, e.g., over 2$\times$ faster than the MAE baseline.
-Besides, it outperforms RVSA and others when scaled to ViT-L.
-This demonstrates that, with our SelectiveMAE training method on the RS-4M dataset, the model can learn strong feature representations and scale efficiently.
 Horizontal & Oriented Object Detection.
@@
 The results are summarized in Table 6.
-Our approach, utilizing a ViT-B backbone, demonstrates competitive or superior performance compared to other methods with a Swin-B backbone, such as RingMo [37], SatLas [83], and GFM [25].
-This is particularly noteworthy as hierarchical transformers generally excel at generating multiscale features essential for object detection.
-Moreover, our SelectiveMAE with a ViT-B backbone outperforms SatMAE [30] and ScaleMAE [33], which use a larger ViT-L backbone, and matches the advanced RVSA [18, 34] that employs specialized window attention.
-When using the larger ViT-L backbone, SelectiveMAE shows significantly enhanced performance across both detection datasets, underscoring the excellent scalability of our method.
+Our approach, utilizing a ViT-B backbone, demonstrates competitive or superior performance compared to other methods with a Swin-B backbone, such as RingMo [37].
+When using the larger ViT-L backbone, SelectiveMAE shows enhanced performance across both detection datasets, underscoring the excellent scalability of our method.
 Semantic Segmentation.
@@
 5 Conclusion
-In this paper, we introduce RS-4M, a large-scale optical remote sensing dataset for unsupervised learning.
-Unlike previous datasets in the RS field, RS-4M offers a significantly larger and more diverse image set with fine-grained details relevant to downstream tasks.
-Benchmarking representative MIM methods on RS-4M highlights its advantages in these tasks.
-To address the computational burden of MIM training on large-scale RS datasets, we propose SelectiveMAE, an efficient MIM method that dynamically encodes and reconstructs patch tokens based on their semantic richness.
-SelectiveMAE significantly accelerates training, demonstrating that only 40% of RS image patches are needed to train a comparable MIM model.
-Experiments show that SelectiveMAE not only achieves over 2×2$\times$2 × speedup in pre-training but also enhances performance in various downstream tasks.
-Limitations and Discussion.
-Although the RS-4M dataset surpasses existing RS datasets by more than four times, its size can be further increased by collecting and filtering additional RS data.
-Additionally, its diversity can be enhanced by incorporating data from various types of sensors.
-    A: The dataset was created by:
-    - Fengxiang Wang (National University of Defense Technology),
-    - Hongzhen Wang (Tsinghua University),
-    - Di Wang (Wuhan University),
-    - Zonghao Guo (University of Chinese Academic of Sciences),
-    - Zhenyu Zhong (Nankai University),
-    - Long Lan (National University of Defense Technology),
-    - Jing Zhang (The University of Sydney),
-    - Zhiyuan Liu (Tsinghua University),
-    - Maosong Sun (Tsinghua University).
-  3. 3.
-    “Who funded the creation of the dataset?”
-    A: The dataset creation was funded by the affiliations of the authors involved in this work.
-7.5.2 Composition
-Most of the questions in this section are intended to provide dataset consumers with the information they need to make informed decisions about using the dataset for their chosen tasks.
-Some of the questions are designed to elicit information about compliance with the EU’s General Data Protection Regulation (GDPR) or comparable regulations in other jurisdictions.
-Questions that apply only to datasets that relate to people are grouped together at the end of the section.
-We recommend taking a broad interpretation of whether a dataset relates to people.
-For example, any dataset containing text that was written by people relates to people.
-  1. 1.
-    “What do the instances that comprise our datasets represent (e.g., documents, photos, people, countries)?”
-    A: The dataset primarily comprises visible light remote sensing images captured by satellites. All datasets utilized in RS-4M are publicly accessible and nonprofit.
-  2. 2.
-    “How many instances are there in total (of each type, if appropriate)?”
-    A: RS-4M contains 4 million remote sensing image instances captured by satellites.
-  3. 3.
-    “Does the dataset contain all possible instances or is it a sample (not necessarily random) of instances from a larger set?”
-    A: Yes, our dataset contains all possible instances that have been collected so far.
-  4. 4.
-    “Is there a label or target associated with each instance?”
-    A: No, our dataset is intended for self-supervised learning. Therefore, each instance is an individual remote sensing image and does not contain annotations.
-  5. 5.
-    “Is any information missing from individual instances?”
-    A: No.
-  6. 6.
-    “Are relationships between individual instances made explicit (e.g., users’ movie ratings, social network links)?”
-    A: Yes, the relationship between individual instances is explicit.
-  7. 7.
-    “Are there recommended data splits (e.g., training, development/validation, testing)?”
-    A: Yes, the entire dataset is intended for self-supervised methods, and we recommend using the whole dataset for self-supervised learning research.
-  8. 8.
-    “Is the dataset self-contained, or does it link to or otherwise rely on external resources (e.g., websites, tweets, other datasets)?”
-    A: Yes, our dataset relies on many publicly available remote sensing datasets, which we have detailed in the main text.
-  9. 9.
-    “Does the dataset contain data that might be considered confidential (e.g., data that is protected by legal privilege or by doctor–patient confidentiality, data that includes the content of individuals’ non-public communications)?”
-    A: No, all data are clearly licensed.
-  10. 10.
-    “Does the dataset contain data that, if viewed directly, might be offensive, insulting, threatening, or might otherwise cause anxiety?”
-    A: No.
-7.5.3 Collection Process
-In addition to the goals outlined in the previous section, the questions in this section are designed to elicit information that may help researchers and practitioners create alternative datasets with similar characteristics.
-Again, questions that apply only to datasets that relate to people are grouped together at the end of the section.
-  1. 1.
-    “How was the data associated with each instance acquired?”
-    A: Please refer to the details listed in the main text Sec. 3.
-  2. 2.
-    “What mechanisms or procedures were used to collect the data (e.g., hardware apparatuses or sensors, manual human curation, software programs, software APIs)?”
-    A: Please refer to the details listed in the main text Sec. 3.
-  3. 3.
-    “If the dataset is a sample from a larger set, what was the sampling strategy (e.g., deterministic, probabilistic with specific sampling probabilities)?”
-    A: Please refer to the details listed in the main text Sec. 3.
-7.5.4 Preprocessing, Cleaning, and Labeling
-The questions in this section are intended to provide dataset consumers with the information they need to determine whether the “raw” data has been processed in ways that are compatible with their chosen tasks.
-For example, text that has been converted into a “bag-of-words" is not suitable for tasks involving word order.
-  1. 1.
-    “Was any preprocessing/cleaning/labeling of the data done (e.g., discretization or bucketing, tokenization, part-of-speech tagging, SIFT feature extraction, removal of instances, processing of missing values)?”
-    A: Yes, we preprocessed and cleaned data in our dataset.
-  2. 2.
-    “Was the ‘raw’ data saved in addition to the preprocessed/cleaned/labeled data (e.g., to support unanticipated future uses)?”
-    A: Yes, raw data is accessible.
-  3. 3.
-    “Is the software that was used to preprocess/clean/label the data available?”
-    A: Yes, the necessary software used to preprocess and clean the data is publicly available.
-7.5.5 Uses
-The questions in this section are intended to encourage dataset creators to reflect on tasks for which the dataset should and should not be used.
-By explicitly highlighting these tasks, dataset creators can help dataset consumers make informed decisions, thereby avoiding potential risks or harms.
-  1. 1.
-    “Has the dataset been used for any tasks already?”
-    A: No.
-  2. 2.
-    “Is there a repository that links to any or all papers or systems that use the dataset?”
-    A: Yes, we provide such links in our GitHub repository.
-  3. 3.
-    “What (other) tasks could the dataset be used for?”
-    A: The dataset could be used for training the remote sensing foundation models with the self-supervised learning method.
-  4. 4.
-    “Is there anything about the composition of the dataset or the way it was collected and preprocessed/cleaned/labeled that might impact future uses?”
-    A: N/A.
-  5. 5.
-    “Are there tasks for which the dataset should not be used?”
-    A: N/A.
-7.5.6 Distribution
-Dataset creators should provide answers to these questions prior to distributing the dataset either internally within the entity on behalf of which the dataset was created or externally to third parties.
-  1. 1.
-    “Will the dataset be distributed to third parties outside of the entity (e.g., company, institution, organization) on behalf of which the dataset was created?”
-    A: No.
-  2. 2.
-    “How will the dataset be distributed (e.g., tarball on website, API, GitHub)?”
-    A: Very likely to be distributed by website, API, and GitHub repository.
-  3. 3.
-    “When will the dataset be distributed?”
-    A: The datasets are publicly accessible.
-  4. 4.
-    “Will the dataset be distributed under a copyright or other intellectual property (IP) license, and/or under applicable terms of use (ToU)?”
-    A: Yes, the dataset is under the Creative Commons Attribution-NonCommercial-ShareAlike 4.0 International License.
-  5. 5.
-    “Have any third parties imposed IP-based or other restrictions on the data associated with the instances?”
-    A: No.
-  6. 6.
-    “Do any export controls or other regulatory restrictions apply to the dataset or to individual instances?”
-    A: No.
-7.5.7 Maintenance
-As with the questions in the previous section, dataset creators should provide answers to these questions prior to distributing the dataset.
-The questions in this section are intended to encourage dataset creators to plan for dataset maintenance and communicate this plan to dataset consumers.
-  1. 1.
-    “Who will be supporting/hosting/maintaining the dataset?”
-    A: The authors of this work serve to support, host, and maintain the datasets.
-  2. 2.
-    “How can the owner/curator/manager of the dataset be contacted (e.g., email address)?”
-    A: The curators can be contacted via the email addresses listed on our webpage [^4] https://github.com/Fengxiang23/SelectiveMAE..
-  3. 3.
-    “Is there an erratum?”
-    A: There is no explicit erratum; updates and known errors will be specified in future versions.
-  4. 4.
-    “Will the dataset be updated (e.g., to correct labeling errors, add new instances, delete instances)?”
-    A: Yes, for the current version. Future updates (if any) will be posted on the dataset website.
-  5. 5.
-    “Will older versions of the dataset continue to be supported/hosted/maintained?”
-    A: Yes. This is the first version of the release; future updates will be posted and older versions will be replaced.
-  6. 6.
-    “If others want to extend/augment/build on/contribute to the dataset, is there a mechanism for them to do so?”
-    A: Yes, we provide detailed instructions for future extensions.
-7.6 Limitation and Potential Societal Impact
-In this section, we elaborate on the limitations and potential societal impact of this work.
-7.6.1 Potential Limitations
-While RS-4M provides a comprehensive benchmark for training the remote sensing foundation models with self-supervised learning methods, there are several limitations to consider:
-  • Scope of Sensors: Although our benchmark includes 4 million visible light remote sensing images, it may not cover all possible real-world scenarios.
-There could be additional sensor data, like multispectral data that were not included in this study, potentially limiting the generalizability of our findings.
-  • Model and Dataset Diversity: While our dataset is primarily focused on adapting to downstream tasks like detection and segmentation, there is undeniably less data available for these tasks compared to general scene classification.
-In the future, we should collect more data that is better suited for various downstream tasks.
-  • Computation and Resource Requirements: Pre-training on the extensive RS-4M dataset, comprising 4 million images, demands substantial computational resources, despite our introduction of the efficient SelectiveMAE method.
-This may limit access to the benchmark for research groups lacking ample computational power.
-7.6.2 Potential Negative Societal Impact
-While the development of remote sensing foundation models with self-supervised learning methods has the potential to significantly advance remote sensing downstream tasks, there are potential negative societal impacts that must be considered:
-  • Safety Risks: Our benchmark aims to improve remote sensing foundation models, but relying too heavily on these models may breed overconfidence in autonomous systems.
-It’s crucial to deploy these systems with adequate safety measures and human oversight to uphold public safety.
-  • Environmental Impact: Training and evaluating remote sensing foundation models demand substantial computational resources, leading to a notable environmental footprint.
-Encouraging the adoption of energy-efficient algorithms (i.e., the proposed SelectiveMAE) and sustainable computing practices is crucial to reduce the environmental impact of this research.
-  • Bias and Fairness: The performance of foundational remote sensing models can vary across different environments and conditions, potentially introducing biases in downstream tasks within the remote sensing domain.
-It is essential to train and evaluate these models on diverse datasets to ensure fairness and avoid discriminatory outcomes.
+In this paper, we introduce a new pre-training pipeline for RS models, featuring the creation of a large-scale RS dataset and an efficient MIM approach.
+Built upon it, we first curated OpticalRS-4M, a large-scale optical remote sensing dataset for unsupervised learning.
+Unlike previous RS datasets, OpticalRS-4M offers a larger and more diverse image set with fine-grained details relevant to downstream tasks.
+Benchmarking representative MIM methods on OpticalRS-4M highlights its advantages in these tasks.
+Then, we present SelectiveMAE to reduce the computational overhead of MIM training on large-scale RS datasets.
+This efficient MIM method dynamically encodes and reconstructs tokens based on their semantic richness.
+SelectiveMAE significantly accelerates training, demonstrating that only 40% RS image patches are needed to train a comparable MIM model.
+Extensive experiments demonstrate that OpticalRS-4M significantly improves classification, detection, and segmentation performance, while SelectiveMAE achieves over 2×2$\times$2 × speedup, highlighting the effectiveness and scalability of our pipeline in developing RS foundational models.

### v4

@@
-OpticalRS-4M: Scaling Efficient Masked Autoencoder Learning on Large Remote Sensing Dataset
+Scaling Efficient Masked Image Modeling on Large Remote Sensing Dataset
 Abstract
@@
 To address these issues, we present a new pre-training pipeline for RS models, featuring the creation of a large-scale RS dataset and an efficient MIM approach.
-We curated a high-quality dataset named OpticalRS-4M by collecting publicly available RS datasets and processing them through exclusion, slicing, and deduplication.
-OpticalRS-4M comprises 4 million optical images covering various RS tasks, such as object detection and pixel segmentation.
+We curated a high-quality dataset named OpticalRS-13M by collecting publicly available RS datasets and processing them through exclusion, slicing, and deduplication.
+OpticalRS-13M comprises 13 million optical images covering various RS tasks, such as object detection and pixel segmentation.
 To enhance efficiency, we propose SelectiveMAE, a pre-training method that dynamically encodes and reconstructs semantically rich patch tokens, thereby reducing the inefficiencies of traditional MIM models caused by redundant background pixels in RS images.
-Extensive experiments demonstrate that OpticalRS-4M significantly improves classification, detection, and segmentation performance, while SelectiveMAE increases training efficiency over 2$\times$ times.
+Extensive experiments demonstrate that OpticalRS-13M significantly improves classification, detection, and segmentation performance, while SelectiveMAE increases training efficiency over 2$\times$ times.
 This highlights the effectiveness and scalability of our pipeline in developing RS foundational models.
-The dataset, source code, and trained models will be released at OpticalRS-4M.
 1 Introduction
+Figure 1: Advantages of SelectiveMAE on finetuning performance and pre-training efficiency.
+The accuracy is evaluated on the DIOR detection dataset [41], while the training efficiency is measured on NVIDIA A100 GPUs, with different versions of ViT [19].
 Over the past decade, advancements in remote sensing (RS) technology and data acquisition have improved applications in ecosystem monitoring [54], natural disaster management [6], among others [32, 74].
@@
 However, each of these downstream tasks often requires substantial computational resources to learn task-specific feature representations and develop specialized models.
+Figure 2: Advantages of OpticalRS-13M in data volumn and diversity.
+(a) The volumn of OpticalRS-13M surpasses at least 4$\times$ compared to existing RS datasets.
+(b) OpticalRS-13M includes significantly more sample categories than others, spanning both object and event types.
+(c) OpticalRS-13M provides more abundant feature representations compared to other large-scale RS datasets.
+Here, 1,000 samples are separately selected from different datasets, and features are extracted using a pre-trained ViT-B.
 Thanks to advances in self-supervised learning methods, such as Masked Image Modeling (MIM) techniques [28, 89], the pre-training of visual foundation models has seen remarkable improvements [90, 94, 91, 95, 73, 77].
@@
 This limitation restricts the generalization of the learned representations.
-To address these challenges, we propose a new pipeline to collect, create, and efficiently process a large optical RS dataset.
+To address these challenges, taking an example of visible light RS imagery, we propose a new pipeline to collect, create, and efficiently process a large RS dataset.
 Firstly, we reviewed publicly available remote sensing datasets from the past decade and selected them based on the DiRS (Diversity, Richness, and Scalability) principle [48].
@@
 Therefore, we also apply exclusion, slicing, and deduplication processes to further improve data quality.
-Correspondingly, we obtain a large-scale RS dataset called OpticalRS-4M comprising 4 million optical images, which is designed to fully leverage the representation learning capabilities of MIM methods in RS applications.
-OpticalRS-4M exceeds previous RS datasets [24, 21, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 18, 37, 38], being at least four times larger.
-Moreover, OpticalRS-4M encompasses a wide range of diverse RS scenarios encountered in downstream tasks such as object-level detection and pixel-level segmentation (Fig.
+Correspondingly, we obtain a large-scale RS dataset called OpticalRS-13M comprising 13 million images, which is designed to fully leverage the representation learning capabilities of MIM methods in RS applications.
+OpticalRS-13M exceeds previous RS datasets [1, 50, 51, 33, 71, 65, 10, 17, 92, 35, 42, 56, 72, 64, 53, 70, 63, 52], being at least four times larger (Fig.
 2 left).
-Despite substantial efforts in training RSFM using MIM methods, such as leveraging general image knowledge [25, 26, 28], expanding parameter scales [29], integrating spatio-temporal information [30, 31, 32], and learning multi-scale features [33, 35, 36], the computational burden and slow convergence when employing MIM training on large-scale RS datasets cannot be ignored.
+Moreover, OpticalRS-13M encompasses a wide range of diverse RS scenarios encountered in downstream tasks such as object-level detection and pixel-level segmentation (Fig.
+2 right).
+Despite substantial efforts in training RSFM using MIM methods [51, 18, 17, 56], the computational burden and slow convergence when employing MIM training on large-scale RS datasets cannot be ignored.
 Specifically, pre-training on 1 million RS samples requires 107 hours for the ViT-B [19] backbone on 8 Nvidia A100 GPUs [70].
-This issue becomes even more pronounced when training on larger datasets, e.g., OpticalRS-4M.
+This issue becomes even more pronounced when training on larger datasets, e.g., OpticalRS-13M.
 In natural scene analysis, this issue has led to numerous studies [25, 75, 40, 11, 23, 20] aimed at improving MIM training efficiency.
@@
 This raises two key questions about how to efficiently conduct MIM training in the RS field: 1) Is it necessary to reconstruct all the redundant background patches during the MIM decoding process?
-2) Is there a feasible way to encode fewer image patches ($\bm{e.g.,\leq}$25%) to accelerate the convergence of MIM training?
+2) Is there a feasible way to encode fewer image patches (e.g., $\leq$25%) to accelerate the convergence of MIM training?
 To address the first question, a measure-based selection process is needed to identify the appropriate patches for reconstruction.
@@
 Then, it selects a subset of patch tokens (e.g., $\leq$50%) with higher HOG values for feature encoding (e.g., $\leq$15%) and pixel reconstruction (e.g., $\leq$35%).
-However, using an extremely low ratio of visible patches during MIM training can lead to gradient explosion or vanishing.
+However, using an extremely low ratio of visible patches during MIM training can lead to gradient explosion.
 To mitigate this, we designed a Progressive Semantic Token Selection (PSTS) module, which dynamically selects semantically relevant patch tokens during the entire training phase.
@@
 As training advances, SelectiveMAE shifts to reconstructing high-semantic tokens from encoded lower-semantic ones to capture complementary semantic dependencies.
-This analogical-to-complementary strategy allows SelectiveMAE to efficiently and progressively learn robust representations of RS images while accelerating MIM convergence.
+This analogical-to-complementary strategy allows SelectiveMAE to efficiently and progressively learn robust representations of RS images while accelerating MIM convergence (Fig.
+1).
 Our experiments reveal that 40% of RS image patches are sufficient to train a comparable model, offering new insights into MIM training on RS images.
-The contributions of this paper are summarized as follows:
+In summary, our main contributions are as follows:
 1) We introduce a new pipeline to collect, create, and efficiently process a large optical RS dataset for developing RS foundation models.
-2) Using this pipeline, we create the OpticalRS-4M dataset, which is a large-scale RS dataset comprising 4 million optical images with diversified coverage scenarios.
+2) Using this pipeline, we create the OpticalRS-13M dataset, which is a large-scale RS dataset comprising 13 million optical images with diversified coverage scenarios.
 3) We introduce SelectiveMAE, an efficient MIM method tailored for RS image pre-training.
@@
 Experiment results suggest the effectiveness and scalability of the proposed pipeline.
-OpticalRS-4M significantly enhances the performance of RS foundation models in downstream tasks, while SelectiveMAE achieves over 2×2$\times$2 × speedup in pre-training compared to MAE [10].
+OpticalRS-13M significantly enhances the performance of RS foundation models in downstream tasks, while SelectiveMAE achieves over 2×2$\times$2 × speedup in pre-training compared to MAE [28].
 2 Related Work
@@
 Some works combine various sensor data to create extensive datasets.
-SEN12MS [53] with 180,662 triplets of dual-pol synthetic aperture radar (SAR) and multi-spectral Sentinel-2 image patches.
-SSL4EO-S12 [54] dataset uses publicly available SAR/optical satellite data collected by the European Space Agency’s Sentinel mission.
-SSL4EO-L [55] consists entirely of multispectral data.
-SatlasPretrain [56] is constructed from data sources such as Sentinel-2, Sentinel-1, Landsat 8/9, and 0.5 - 2 meter/pixel aerial imagery.
-While it includes millions of images, the majority are multispectral and SAR, with only a small fraction being RGB.
+The SEN12MS dataset [59] consists of 180,662 triplets, each containing synthetic aperture radar (SAR) and multispectral Sentinel-2 image patches, and MODIS land cover maps.
+While SSL4EO-S12 [79] , SSL4EO-L [61] and SatlasPretrain [5] contain millions of images, the major data type in these datasets are multispectral and SAR, with only a small fraction being RGB.
 There are also some large-scale RS visual-language datasets such as Skyscript [80] and RS5M [96], but they primarily focus on multimodal tasks.
-Currently, some datasets focus solely on optical RS images.
-MillionAID [20] offers a million-scale dataset for RS optical image classification, while SeCo [22] and CACo [21] provide nearly a million images of the same location over different times.
+Currently, there are also some datasets that focus solely on visible light RS images.
+MillionAID [48], SeCo [49] and CACo [50] provide nearly a million images of the same location over different times.
 However, these datasets primarily target scene classification and often overlook fine-grained target information, limiting their utility for various downstream tasks.
-To address this gap, with the proposed pipeline, we introduce the OpticalRS-4M dataset, which is larger and more diverse, enhancing performance across multiple downstream tasks.
+To address this gap, with the proposed pipeline, we introduce the OpticalRS-13M dataset, which is larger and more diverse, enhancing performance across multiple downstream tasks.
 Masked Image Modeling.
@@
 MIM learns image representations by reconstructing masked tokens, focusing on various regression targets [29, 4, 82, 97, 47], masking strategies [60, 39], and reconstruction methods [67, 83, 45, 36].
-For instance, MAE [10] demonstrates that predicting pixel values can be as effective as using complex targets.
 A major challenge for MIM is its high computational demand and lengthy pre-training times.
-To mitigate this, some studies use asymmetric encoder-decoder strategies [40, 41], reduce input patches [42, 44] or use a novel Difficulty-Flatten Loss [43].
+To mitigate this, approaches such as asymmetric encoder-decoder strategies [25, 75], reducing input patches [40, 23], or the Difficulty-Flatten Loss [11] have been proposed.
 Additionally, CrossMAE [20] employs cross-attention between masked and visible tokens to enhance efficiency without sacrificing performance.
@@
 Self-supervised learning methods have recently been employed to extract representations from unlabeled RS data.
-Although there are some contrastive self-supervised methods [24, 21, 22, 75], recent advancements have primarily centered around generative self-supervised methods.
-For example, RVSA [18] leverages the vision transformer backbone pre-trained by the vanilla MAE [10] method and introduces computationally efficient rotated varied-size window attention to replace the original attention during fine-tuning.
-In addition, many studies are focusing on improving generative self-supervised algorithms by leveraging general image knowledge [25, 26, 28], scaling up parameter sizes [29], integrating spatio-temporal data [30, 31, 32], handling multi-sensor data [76, 77, 78, 79, 80], and employing multi-scale concepts [33, 35, 36].
+Due to the inefficiency of designing pretext tasks and gathering required data for contrastive self-supervised methods [1, 50, 49, 81], recent advancements have primarily centered around generative self-supervised methods, especially in MIM.
+Specifically, many studies aim to improve generative self-supervised algorithms by leveraging general image knowledge [51, 33, 65], scaling up parameter sizes [10, 18], integrating spatio-temporal information [17, 92, 42], encompassing geometric attributes [70, 46], handling multi-sensor data [34, 24, 27, 8, 31], and employing multi-scale concepts [56, 64, 53].
 However, these methods have not effectively addressed the substantial computational burden associated with self-supervised pre-training in RS.
 In the paper, we propose a new pipeline that can collect, create, and efficiently process large amounts of optical RS data for developing RS foundation models, enhancing the practicality of MIM pre-training on large-scale datasets.
-3 Methodology
+3 Method
 The proposed pipeline contains two critical components: dataset generation and efficient pre-training, which will be introduced in detail in the following text.
-3.1 Dataset Generation
-Figure 1: Left: The organization of data sources in OpticalRS-4M. Right: The comparison between OpticalRS-4M and other datasets.
+3.1 Dataset Curation
 Recent progress in self-supervised pre-training for RSFMs is limited by the relatively small scale and diversity of existing RS datasets compared to natural scene datasets.
-To overcome this, we first curate a large-scale RS dataset, OpticalRS-4M, through extensive data collection and preprocessing.
-3.1.1 Data Collection
-We began by reviewing publicly available RS datasets from the past decade.
+To overcome this, we curate a large-scale RS dataset, OpticalRS-13M, through extensive data collection and preprocessing.
+Data Collection We began by reviewing publicly available RS datasets from the past decade.
 Recognizing that not all data are suitable for self-supervised pre-training, we adopted specific criteria for data collection, as outlined in DiRS [48]: 1) Diversity: A dataset is diverse if its images capture various typical visual features of relevant scenes or targets and offer complementarity.
-Intra-class Diversity: High diversity within classes ensures comprehensive representation of real-world target distributions.
-To ensure this, we consider the data from different sources of the same scene or target.
-Moreover, in the RS field, the attributes of objects may vary with geographic location and imaging time, so the change detection data is also incorporated.
-Inter-class Diversity: To enable self-supervised pre-training to distinguish between categories effectively, we incorporated as many fine-grained categories as possible, especially those with high semantic overlap.
-This strategy helps the model learn robust feature representations.
-2) Richness: Beyond diversity, richness is also essential.
-We ensure diverse content features and large sample sizes by collecting images under various conditions — weather, seasons, lighting, and so on — introducing variability in translation, viewpoint, object posture, spatial resolution, background, and occlusion.
-Our data collection process actively includes the sample of the same category under different imaging conditions to ensure comprehensive real-world representation, enhancing the model’s representation and generalization capabilities.
-3) Scalability: Given the evolving applications of remote sensing images, scalability is also crucial.
-We constantly enrich the dataset based on scene and target categories across classification, detection, and segmentation tasks.
-In summary, following the guidelines, we constructed a large-scale RS image dataset with diverse coverage scenarios. Fig. 1 left illustrates the included RS datasets collected.
-3.1.2 Data Preprocessing
-Following data collection, challenges such as inconsistent data sources, oversized images, and redundant pixels remained.
+Intra-class Diversity: High within-class diversity ensures comprehensive representation of real-world target distributions.
+We include data from different sources of the same scene or target, incorporating change detection data to capture object variations by geographic location and imaging time.
+Inter-class Diversity: To enable self-supervised pretraining to distinguish categories effectively, we included as many fine-grained categories as possible, especially those with high semantic overlap, helping the model learn robust feature representations.
+2) Richness: Beyond diversity, richness is essential.
+We ensure varied content features and large sample sizes by collecting images under multiple conditions (e.g., weather, seasons, lighting), introducing variability in translation, viewpoint, object posture, spatial resolution, background, and occlusion.
+Sampling across imaging conditions provides a comprehensive real-world representation, enhancing model’s representation and generalization capabilities.
+3) Scalability: Scalability is crucial given the evolving applications of remote sensing images.
+We continuously expand the dataset based on scene and target categories across classification, detection, and segmentation tasks.
+In summary, following the guidelines, we constructed a large-scale RS image dataset with diverse coverage scenarios.
+Data Preprocessing Following data collection, challenges such as inconsistent data sources, oversized images, and redundant pixels remained.
 To address these, we implemented a standardized, scalable data preprocessing workflow consisting of four steps:
-1) We focused exclusively on optical images for this study, excluding multispectral and SAR data, though future updates will incorporate these modalities for multi-modal self-supervised pre-training.
+1) We focused exclusively on visible light images for this study, excluding multispectral and SAR data, though future updates will incorporate these modalities for multi-modal self-supervised pre-training.
 2) To manage the large image sizes, we randomly cropped high-resolution images into smaller slices.
@@
 This ensured that only highly similar images were excluded.
-As a result of these steps, we created the OpticalRS-4M dataset, which contains approximately 4 million high-quality RS images — significantly larger than previous representative RS datasets, as illustrated in Fig.
+As a result of these steps, we created the OpticalRS-13M dataset, which contains approximately 13 million high-quality RS images, significantly larger than previous representative RS datasets, as illustrated in Fig.
+2 left.
+In summary, we followed the DiRS principles for data collection, applying exclusion, slicing, and deduplication to ensure high-quality.
+We’ve screened and processed most small and medium-scale datasets from the past decade in this field.
+Dataset Diversity To demonstrate the broad scope of scenes and target categories covered by our dataset, we conducted a statistical analysis accompanied by a detailed label explanation, as shown in Fig.
 2 right.
-In summary, we followed the DiRS principles for data collection, applying exclusion, slicing, and deduplication to ensure high-quality results.
-We’ve screened and processed most small and medium-scale datasets from the past decade in this field.
-More efforts can be made to extend OpticalRS-4M by including larger datasets like SSL4EO-S12 [54] and SSL4EO-L [55] to enhance dataset capacity.
-3.1.3 Dataset Diversity
-Refer to captionRefer to caption
-Figure 2: Left: Main categories with visual examples from OpticalRS-4M.Right: Full hierarchy of OpticalRS-4M categories.
-To demonstrate the broad scope of scenes and target categories covered by our dataset, we conducted a statistical analysis accompanied by a detailed label explanation, reviewing all the scenarios and targets included in OpticalRS-4M, as shown in Fig.
-2.
-The analysis reveals that OpticalRS-4M comprises 12 main categories, each containing numerous subclasses.
-Unlike existing self-supervised optical remote sensing datasets, OpticalRS-4M not only offers more comprehensive scene and target information but also introduces “Events” as a distinct high-level category.
-This category includes labels such as “Concert”, “Parade/Protest”, “Religious Activity”, “Fire”, “Flood”, “Landslide”, “Post-Earthquake”, and “Traffic Collision”.
-This expanded information enhances the versatility of OpticalRS-4M for a wider range of downstream tasks.
+The analysis reveals that OpticalRS-13M comprises 12 main categories, each containing numerous subclasses.
+Unlike existing self-supervised optical remote sensing datasets, OpticalRS-13M not only offers more comprehensive scene and target information but also introduces “Events” as a distinct high-level category.
+This category includes labels such as “Fire”, “Flood”, “Landslide”, “Post-Earthquake”, and so on.
+This expanded information enhances the versatility of OpticalRS-13M for a wider range of downstream tasks.
 3.2 Efficient Pre-training
-After obtaining a large-scale RS dataset, the next step is to conduct the self-supervised pre-training.
+After obtaining a large-scale RS dataset, the next step is conduct the self-supervised pre-training.
 However, as the capacity of the dataset grows, vast computational and time costs are required for existing MIM approaches used in the RS community, e.g., MAE [28].
-To address this issue, our pipeline introduces an efficient MIM method, i.e., SelectiveMAE.
+To address this issue, our pipeline introduces an efficient MIM method, SelectiveMAE.
 In this section, we first review the preliminaries of MAE, then we present the details of SelectiveMAE.
-3.2.1 Masked Autoencoders Preliminaries
+3.2.1 Preliminaries of MAE
+Figure 3: Overview of SelectiveMAE. It inputs fewer visible patches and reconstructs only partial patches to accelerate training efficiency.
 1) Masking.
@@
 It linearly projects the patches, adds positional embeddings, and processes them through a series of transformer blocks.
-By operating on a smaller subset of patches, the encoder enables the training of large models with reduced computational and memory requirements.
-3) MAE Decoder.The encoded tokens and masked tokens are fed into the decoder, which comprises transformer blocks with self-attention layers.
+By operating on a smaller subset of patches, the encoder enables training of large models with reduced computational and memory requirements.
+3) MAE Decoder.
+The encoded tokens and masked tokens are fed into the decoder, which comprises transformer blocks with self-attention layers.
 The masked tokens are shared, learnable tensors enhanced with positional embeddings.
@@
 The loss function computes the mean squared error (MSE) between the reconstructed targets and original patches.
-RS optical images typically contain many redundant background pixels, which aligns well with MAE’s masking strategy.
-In MAE, the encoder processes only 25% of the patches, reducing computational load by avoiding processing the entire image.
-Building on MAE, we aim to leverage the redundancy in RS images to accelerate training.
-Specifically, we address two questions: 1) Is it necessary to reconstruct all the masked patches given the redundancy in RS images?
+However, applying MAE for self-supervised pre-training results in considerable training costs in terms of time complexity, especially when working with large RS datasets.
+To address this, we observe that RS optical images often contain redundant information, which could be omitted during training to improve pre-training efficiency.
+Specifically, we address two issues: 1) Is it necessary to reconstruct all the masked patches given the redundancy in RS images?
 2) Can the visible patches input to the MAE encoder be further compressed to enhance acceleration?
-Figure 3: Overview of SelectiveMAE. The vanilla version of SelectiveMAE.
 3.2.2 Partial Reconstruction
@@
 The decoder uses a lightweight design based on cross-attention following CrossMAE [20].
-Experimental results in the Appendix show that this partial reconstruction strategy significantly increases the training throughput without affecting the learned representations.
+Experimental results in supplementary material show that this partial reconstruction strategy significantly increases the training throughput without affecting the learned representations.
 3.2.3 Progressive Semantic Token Selection
+Figure 4: Effectiveness of PSTS. Left: Using only 40% of patches for encoding and reconstruction often leads to gradient explosions. Right: The training loss after adopting PSTS.
 For the second question, we initially tried a naive approach by increasing the masking ratio to 85%, meaning only 15% of the patches in each RS image were input to the encoder while keeping a 25% reconstruction ratio as proposed in Sec.
 3.2.2.
-However, during training, this setup often led to issues like gradient explosions or loss divergence (see the Appendix).
-We speculate that, although RS images have background redundancy, using such a small portion of patches challenges MAE training.
-How can we achieve acceleration at a high masking ratio while ensuring MAE convergence?
-Curriculum learning [82, 83, 84] may offer a solution.
-However, it typically focuses on selecting whole image samples, whereas MIM pre-training masks large portions of image pixels, emphasizing patch-level selection.
-Therefore, it’s essential to distinguish between “easy” and “hard” patches in MIM pre-training.
-To this end, we introduce the Progressive Semantic Token Selection (PSTS) module for patch selection.
-In this module, depicted in Fig.
-3, we begin by selecting a limited number of patches and then select additional patches based on them in the training epoch, dynamically transitioning from easily learned, semantically similar patches to more challenging, complementary ones.
-For initialization, we employ a HOG selection strategy to choose the initial patch set.
+However, during training, this setup often led to issues like gradient explosions or loss divergence, as shown in Fig.
+4 left.
+The figure shows that using an extremely low portion of patches for encoding and reconstruction caused unstable training due to gradient explosions.
+How can we achieve acceleration at a high masking ratio (e.g., 85%) while ensuring MAE completes the pre-training task for RS images?
+Inspired by the curriculum learning  [78, 7, 26], which follows the principle of learning from easy to hard, we introduce the Progressive Semantic Token Selection (PSTS) module for patch selection, as depicted in Fig.
+3.
+In this module, we begin by selecting a limited number of patches and then select additional patches based on them in the training epoch, dynamically transitioning from easily learned, semantically similar patches to more challenging, complementary ones.
+For initialization, we employ a HOG selection strategy to choose the initial patch set from $S^{N}=\left\{x_{i}^{p}\right\}_{i=1}^{N}$, with a proportion $s\in[0,(1-m)/2]$.
+We define the initial set of selected token as:
+[…]
+We select $\lfloor\ s\times N\rfloor$ tokens with the maximum HOG feature values (i.e., ${top}_{\lfloor\ s\times N\rfloor}$) from the original token set to form the initial token set.
 This simple yet effective strategy ensures that semantically rich tokens are selected.
-Following token initialization, we aim to incrementally increase the number of tokens to facilitate a training progression from easy to difficult, while adhering to the final masking ratio for selected tokens.
-This process is outlined in the Appendix.
-Given the initial token sets’ high HOG feature values, nearby tokens selected by PSTS also exhibit high HOG feature values.
-Recalling our partial reconstruction approach in Sec.
-3.2.2, we filter the remaining unselected token set to include those with high HOG values as reconstruction targets.
-Consequently, when selecting nearby tokens, the tokens input for encoding and those chosen for reconstruction are closer, thereby simplifying the training process.
-Conversely, if a token’s feature distribution deviates from others in the set, indicating it is farther away, we consider such patches more challenging to learn.
-Details of the process are presented in the Appendix.
+After token initialization, we incrementally increase the number of tokens to guide training from easier to more challenging examples, while maintaining the final masking ratio for selected tokens, as outlined in Algorithm 1.
+Given the high HOG feature values of the initial token sets, nearby tokens selected by PSTS also exhibit high HOG values.
+In partial reconstruction method in Section 3.2.2, we filter unselected tokens to retain those with high HOG values as reconstruction targets.
+Thus, selecting nearby tokens brings encoding tokens and reconstruction targets closer, streamlining training.
+In contrast, if a token’s feature distribution differs significantly from others, we consider these more challenging patches to learn.
+Algorithm 1 Progressive Semantic Token Selection
+1:Number of training epochs $T$ and total training stages $N_{g}$ (i.e., $T/N_{g}$ epochs for each stage), masking rate $m$, input dataset $\mathcal{X}$
+2:Obtain the selected tokens set ${S^{K}}$ and update ${S^{U}}$ in each epoch
+3:for $t\leftarrow 1$ to $T$ do
+4:    Sample data sample from $\mathcal{X}$, feed-forwarded through the embedding to obtain the output token set ${S}_{N}$
+5:    $s\leftarrow\frac{1}{2}(1-m)$
+6:    Obtain ${S}^{I}$ via Eq. (2) and initialize ${S^{U}}$
+7:    Obtain the current training stage $\zeta=\lceil N_{g}*t/T\rceil$
+[…]
+9:    Obtain $S^{*}$ and update $S^{K},S^{U}$ via Eq. (5) and Eq. (6)
+10:end for
+Specifically, we select tokens from $S^{U}$ based on $S^{I}$.
+First, we use $\mathbb{S^{I}}\in\mathbb{R}^{|S^{K}|\times d}$ and $\mathbb{S^{U}}\in\mathbb{R}^{|S^{U}|\times d}$ to denote the matrix representation of the initial token set $S^{I}$ and the unselected token set $S^{U}$, where ||$\cdot$|| ⋅ | represents the number of tokens and $d$ the feature dimension after the embedding layer.
+We use Cosine Distance to measure the dissimilarity between the tokens in these two sets:
+$\small\mathcal{D}(S^{U},S^{I})=\mathbf{1}-\cos\left\langle S^{U},S^{I}\right\rangle=\mathbf{1}-\frac{\mathbb{S^{U}}\mathbb{S^{I}}^{\mathrm{T}}}{\left|\kern-0.96873pt\left|\mathbb{S^{U}}\right|\kern-0.96873pt\right|\cdot\left|\kern-0.96873pt\left|\mathbb{S^{I}}\right|\kern-0.96873pt\right|},$ (3)
+where $\mathbf{1}$ is an all-one matrix.
+$\mathcal{D}(S^{U},S^{I})\in\mathbb{R}^{|S^{U}|\times|S^{I}|}$ represents the pairwise distances between tokens in $S^{U}$ and $S^{I}$.
+Next, we define the distance between the tokens in $S^{U}$ to the initial token set $S^{K}$ based on the selection criteria in each training stage as follows:
+[…]
+where $i\in\{1,\cdots,|S^{U}|\}$, $j\in\{1,\cdots,|S^{I}|\}$, and $\zeta$ represents the training stage depending on the number of epochs.
+Finally, we sample $\lfloor N\times(1-m-s)\rfloor$ tokens from ${S^{U}}$ and add them with ${S^{I}}$ to form ${S^{K}}$, which can be formulated as follows:
+[…]
+$\small{S^{K}}={S^{I}}\cup\mit{S}^{*},\quad{S^{U}}={S^{U}}\setminus\mit{S}^{*},$ (6)
+where $\mit{S}^{*}$ represents the selected token from ${S^{U}}$.
+The operations in Eq.
+(5) and Eq.
+(6) are performed in each training epoch.
+This process is summarized in Algorithm 1.
+Fig.
+5 further provides an example to illustrate which patches should be selected at different stages, helping to facilitate readers’ understanding.
+Figure 5: Illustration of the patch selection process in PSTS.
 4 Experiment
+Table 1: Performance comparison of using the MillionAID and OpticalRS-13M datasets separately with the same data throughput, evaluated by MAE [28] and ViT-B [19].
+Dataset Method Image Number Epoch Scene Object Semantic
+Classification Detection Segmentation
+AID [86] RESISC-45 [13] DIOR [41] DIOR-R [15] LoveDA [76] SpaceNetv1 [68]
+OA (TR=20% / 50%) OA (TR=10% / 20%) $\mathrm{mAP}_{50}$ $\mathrm{mAP}_{50}$ mIoU mF1
+MillionAID [48] MAE 1 million 800 94.92/97.38 89.20/93.60 71.80 62.33 51.24 79.24
+OpticalRS-13M MAE 1 million 800 96.26/97.98 91.53/93.88 75.90 66.46 53.10 79.35
+OpticalRS-13M MAE 2 million 400 96.64/98.10 91.80/94.31 73.90 65.95 52.86 79.37
+OpticalRS-13M MAE 3 million 267 96.67/98.18 92.24/94.41 75.40 67.07 52.39 79.37
+OpticalRS-13M MAE 4 million 200 96.10/98.03 92.38/94.30 74.70 66.26 52.75 79.23
+OpticalRS-13M MAE 8 million 100 96.58/98.26 91.83/93.99 75.20 66.99 52.56 79.43
+OpticalRS-13M MAE 13 million 67 96.28/98.06 91.41/93.60 74.50 66.36 52.63 79.39
 In this section, we first discuss the effectiveness of the dataset and acceleration strategy in the pipeline through ablation experiments.
@@
 Additionally, we performed comprehensive comparative experiments to evaluate the pipeline’s performance on downstream tasks.
-Pre-training Setup.
-Our pre-training experiment setup largely follows MAE [10].
-The masking ratio is 85%, and the reconstruction ratio is 25%.
-For patch selection, the adjustment threshold is set at 300 epochs, with the last 200 epochs using random distance to enhance robustness.
+All experiments are performed by fine-tuning pre-trained models on downstream tasks, including RS scene classification, object detection, and semantic segmentation.
+The pre-training and fine-tuning settings of our methods can be found in the supplementary material.
+Table 2: The performance comparison between SelectiveMAE and the baseline method MAE [28] with the ViT-B backbone.
+The training time is measured on NVIDIA A100 GPUs.
+To accelerate validation, we select 4 million images from OpticalRS-13M for pre-training.
+Method Image Number Epoch Total Training Scene Object Semantic
+Classification Detection Segmentation
+Time (h) AID [86] RESISC-45 [13] DIOR [41] DIOR-R [15] LoveDA [76] SpaceNetv1 [68]
+OA (TR=20% / 50%) OA (TR=10% / 20%) $\mathrm{mAP}_{50}$ $\mathrm{mAP}_{50}$ mIoU mF1
+MAE 4 million 800 177 96.88/98.22 92.44/94.43 75.40 67.35 52.80 79.41
+SelectiveMAE 4 million 800 86 (2.1×2.1$\times$2.1 ×) 96.78/98.12 93.35/94.58 75.70 67.78 53.05 79.50
 4.1 Ablation Study
-To validate the pipeline’s effectiveness, we conducted separate evaluations of both the OpticalRS-4M dataset generated by the pipeline and the accelerated training method, SelectiveMAE.
+To validate the pipeline’s effectiveness, we conducted separate evaluations of both the OpticalRS-13M dataset generated by the pipeline and the accelerated training method, SelectiveMAE.
 Additionally, we assessed the pipeline’s scalability in collecting and processing datasets of varying sizes.
-4.1.1 Superiority of OpticalRS-4M
-Table 1: Comparison of downstream task performance at matched data throughput between MillionAID and OpticalRS-4M, assessed using MAE [10] pre-training and ViT Base models [39].
-Dataset Method Images Number Epoch Sence Object Semantic
-Classification Detection Segmentation
-AID RESISC-45 DIOR DIOR-R LoveDA SpaceNetv1
-OA (TR=20% / 50%) OA (TR=10% / 20%) $\mathrm{mAP}_{50}$ $\mathrm{mAP}_{50}$ mIoU mF1
-MillionAID MAE 1 million 800 94.92/97.38 89.20/93.60 71.80 62.33 51.24 79.24
-OpticalRS-4M MAE 1 million 800 96.26/97.98 91.53/93.88 75.90 66.46 53.10 79.35
-OpticalRS-4M MAE 2 million 400 96.64/98.10 91.80/94.31 73.90 65.95 52.86 79.37
-OpticalRS-4M MAE 3 million 267 96.67/98.18 92.24/94.41 75.40 67.07 52.39 79.37
-OpticalRS-4M MAE 4 million 200 96.10/98.03 92.38/94.30 74.70 66.26 52.75 79.23
-Dataset Diversity.
+More ablation experiment results related to model design are presented in the supplementary material.
+Dataset Diversity of OpticalRS-13M.
 In self-supervised pre-training of RSFMs, the primary goal is to enhance downstream task performance.
-We define dataset diversity as its ability to improve various types of these tasks.
-While datasets like MillionAID focus on classification, our OpticalRS-4M dataset is designed to extend the benefit to detection and segmentation tasks.
+Here, We define dataset diversity as the ability to improve performance across various types of tasks.
+While datasets like MillionAID [48] focus on classification, our OpticalRS-13M dataset is designed to extend the benefit to detection and segmentation tasks.
 We assess this by evaluating the effectiveness of models trained on different datasets with equivalent data throughput.
-For instance, MillionAID, with 1 million samples, is trained using the basic MAE method for 800 epochs as a benchmark.
-To showcase the diversity of OpticalRS-4M, containing 4 million samples, we train it with MAE for just 200 epochs.
-Additionally, we validate OpticalRS-4M’s diversity by progressively reducing its training data — down to 3 million, 2 million, or even 1 million samples — while keeping data throughput constant.
-As shown in Table 1, our experiments demonstrate that even when randomly sampling 1 million images from OpticalRS-4M, it outperforms MillionAID in downstream tasks, particularly in object detection and segmentation.
-This consistent superior performance, even with fewer images, highlights OpticalRS-4M’s greater diversity and effectiveness, emphasizing the benefits of using larger and more varied datasets for unsupervised pre-training.
-The results in Table 1 indicate that the optimal outcomes at equivalent data throughput are achieved with 3 million samples rather than 4 million.
-In most cases, the model pre-trained on 3 million samples outperforms across various downstream tasks.
-Notably, on the DIOR and LoveDA datasets, the best results were obtained with 1 million samples trained for 800 epochs.
-However, as shown in Table 2, when we extended the training on 4 million samples to more epochs, the model achieved the highest average performance across most downstream tasks.
-This suggests that training on 4 million samples for only 200 epochs was insufficient for the MIM method to fully leverage the large dataset, resulting in inferior generalization in downstream tasks.
-Table 2: Performance on downstream tasks with extended training on OpticalRS-4M.
-Dataset Method Images Number Epoch Sence Object Semantic
-Classification Detection Segmentation
-AID RESISC-45 DIOR DIOR-R LoveDA SpaceNetv1
-OA (TR=20% / 50%) OA (TR=10% / 20%) $\mathrm{mAP}_{50}$ $\mathrm{mAP}_{50}$ mIoU mF1
-OpticalRS-4M MAE 3 million 267 96.67/98.18 92.24/94.41 75.40 67.07 52.39 79.37
-OpticalRS-4M MAE 4 million 100 95.12/97.54 91.26/93.85 - - - -
-OpticalRS-4M MAE 4 million 200 96.10/98.03 92.38/94.30 74.70 66.26 52.75 79.23
-OpticalRS-4M MAE 4 million 400 96.36/98.16 92.41/94.03 - - - -
-OpticalRS-4M MAE 4 million 800 96.88/98.22 92.44/94.43 75.40 67.35 52.80 79.41
-Extended training on OpticalRS-4M.
-To maximize the potential of OpticalRS-4M, we extended training using the MAE pre-training method and ViT-B, varying only the number of training epochs while keeping other settings consistent.
-The fine-tuning results, shown in Table 2, reveal that as the number of epochs increases, OpticalRS-4M shows greater advantages in downstream tasks.
-This indicates that extended training is necessary to fully harness OpticalRS-4M’s capabilities.
-However, the gains diminish with additional epochs, suggesting that merely increasing epochs may not suffice.
-To unlock the full potential of OpticalRS-4M, exploring more effective pre-training methods, such as the proposed SelectiveMAE, may be crucial.
-Table 3: Comparison of downstream task performance using SelectiveMAE and MAE [10]. The training time is measured using 8 NVIDIA A100 GPUs.
-Method Dataset Epoch Total Training Sence Object Semantic
-Classification Detection Segmentation
-Hours AID RESISC-45 DIOR DIOR-R LoveDA SpaceNetv1
-OA (TR=20% / 50%) OA (TR=10% / 20%) $\mathrm{mAP}_{50}$ $\mathrm{mAP}_{50}$ mIoU mF1
-MAE OpticalRS-4M 800 177 96.88/98.22 92.44/94.43 75.40 67.35 52.80 79.41
-SelectiveMAE OpticalRS-4M 800 86 (2.1×2.1$\times$2.1 ×) 96.78/98.12 93.35/94.58 75.70 67.78 53.05 79.50
-4.1.2 Effectiveness of SelectiveMAE
-Performance Advantage.
-To further validate the accelerated pre-training strategy of our pipeline, we conducted extensive experiments on OpticalRS-4M using both MAE [10] and SelectiveMAE.
+As shown in Table 1, even with random sampling of 1 million images from OpticalRS-13M, it outperforms MillionAID on downstream tasks, particularly in object detection and segmentation.
+The results in Table 1 also indicate that the optimal outcomes at equivalent data throughput are achieved by different settings other than using all 13 million images.
+This suggests that training on a large number of samples with few epochs was insufficient for the MIM method to fully leverage the large dataset, leading to inferior generalization on downstream tasks.
+Experimental results in the supplementary material show consistent performance improvements across downstream tasks with more training epochs.
+Table 3: The efficiency comparison between SelectiveMAE and the baseline method MAE [28] across different backbones on the MillionAID dataset [48] with 1600 epochs.
+The memory is measured on a single NVIDIA A100 GPU with a batch size of 256.
+Model Backbone Params Total Training Time (h) GPU Memory (MB)
+MAE ViT-B 86M 101 17628
+SelectiveMAE ViT-B 86M 48 (2.1×2.1$\times$2.1 ×) 9570 (1.8×1.8$\times$1.8 ×)
+MAE ViT-L 307M 113 30530
+SelectiveMAE ViT-L 307M 50 (2.3×2.3$\times$2.3 ×) 18790 (1.6×1.6$\times$1.6 ×)
+Performance Advantage of SelectiveMAE.
+To further validate the accelerated pre-training strategy of our pipeline, we conducted extensive experiments on OpticalRS-13M using both MAE [28] and SelectiveMAE.
 The performance of ViT-B on various downstream tasks after pre-training is presented in Table 2.
-The results indicate that after pre-training on OpticalRS-4M, SelectiveMAE not only outperforms MAE but also achieves this at 2.1 times the speed.
-This makes SelectiveMAE particularly well-suited for large-scale datasets, offering significant time savings during training on OpticalRS-4M.
-Table 4: The accelerating performance of SelectiveMAE for different backbones compared to MAE [10]. The training time is measured using 8 NVIDIA A100 GPUs.
-Model Backbone Params Total Training Hours
-MAE ViT-B 86M 101
-SelectiveMAE ViT-B 86M 48 (2.1×2.1$\times$2.1 ×)
-MAE ViT-L 307M 113
-SelectiveMAE ViT-L 307M 50 (2.3×2.3$\times$2.3 ×)
-Pre-training Efficiency.
-To further highlight the efficiency advantage of SelectiveMAE, we evaluate the pretraining time with different backbones on the MillionAID dataset, as shown in Table 4.
-It can be seen that, since only 40% of the image patches (15% for the encoder and 25% for the decoder) are involved in the calculation, our methods present more than doubled acceleration for both base and large ViT models compared to the vanilla MAE baseline, where the acceleration advantage is more significant for larger sized models.
-4.1.3 Scalability of the Pipeline
-Table 5: Training cost (measured using 8 NVIDIA A100 GPUs) of SelectiveMAE and MAE [10] on OpticalRS-4M with different data volumes.
-Method Images Number Epoch Total Training Hours
+The results indicate that after pre-training on OpticalRS-13M, SelectiveMAE not only outperforms MAE but also achieves this at 2.1 times the speed.
+This makes SelectiveMAE particularly well-suited for large-scale datasets, offering significant time savings during training on OpticalRS-13M.
+Efficiency Advantage of SelectiveMAE.
+To further highlight the efficiency advantage of SelectiveMAE, we evaluate the training time and memory footprint during pretraining of different backbones on the MillionAID dataset, as shown in Table 3 (the corresponding accuracies are presented in Table 5).
+It can be seen that, since only 40% of the image patches (15% for the encoder and 25% for the decoder) are involved in the calculation, our methods present more than doubled training acceleration compared to the vanilla MAE baseline, where the acceleration advantage is more significant for larger models.
+Additionally, it is worth noting that SelectiveMAE also reduces memory footprint by using fewer patches.
+Scalability of the Pipeline To evaluate the scalability of our pipeline, we organized datasets containing 1 million and 4 million samples and conducted experiments on each.
+The results in Table 4 show that with 1 million samples, the proposed SelectiveMAE method reduced training time by 29 hours.
+When the dataset was expanded to 4 million samples, the time savings increased to 81 hours.
+These findings demonstrate that our pipeline scales effectively with larger datasets, offering significant pre-training acceleration.
+Table 4: The training cost of SelectiveMAE and the baseline method MAE [28] on OpticalRS-13M with different data volumes. All experiments are conducted on NVIDIA A100 GPUs.
+Method Images Number Epoch Total Training Time (h)
 MAE 1 million 800 54
@@
 SelectiveMAE 4 million 800 86 (-81)
-To evaluate the scalability of our pipeline, we organized datasets containing 1 million and 4 million samples and conducted experiments on each.
-The results, presented in Table 5, show that with 1 million samples, the proposed SelectiveMAE method reduced training time by 29 hours.
-When the dataset was expanded to 4 million samples, the time savings increased to 81 hours.
-These findings demonstrate that our pipeline scales effectively with larger datasets, offering significant pre-training acceleration.
+Table 5: Performance comparison results of different models.
+“TR” represents the ratio of training data to the entire dataset.
+The first and second scores are marked by bold and blue, respectively.
+* means pre-training on 4 million images sampled from OpticalRS-13M.
+Model Backbone Params (M) Scene Object Semantic
+Classification Detection Segmentation
+AID [86] RESISC-45 [13] DIOR [41] DIOR-R [15] LoveDA [76] SpaceNetv1 [68]
+(TR=20%/50%) (TR=10%/20%) mAP50 mAP50 mIoU mF1
+SeCo [49] ResNet-50 26 93.47/95.99 89.64/92.91 - - 43.63 77.09
+GASSL [1] ResNet-50 26 93.55/95.92 90.86/93.06 67.40 65.65 48.76 78.51
+TOV [65] ResNet-50 26 95.16/97.09 90.97/93.79 70.16 66.33 49.70 -
+CACo [50] ResNet-50 26 90.88/95.05 88.28/91.94 66.91 64.10 48.89 77.94
+SatMAE [17] ViT-L 307 95.02/96.94 91.72/94.10 70.89 65.66 - 78.07
+ScaleMAE [56] ViT-L 307 96.44/97.58 92.63/95.04 73.81 66.47 - -
+SSL4EO [79] ViT-S 22 91.06/94.74 87.60/91.27 64.82 61.23 - -
+RingMo [63] Swin-B 88 96.90/98.34 94.25/95.67 75.90 - - -
+SatLas [5] Swin-B 88 94.96/97.38 92.16/94.70 74.10 67.59 - -
+GFM [51] Swin-B 88 95.47/97.09 92.73/94.64 72.84 67.67 - -
+RVSA [70] ViT-B+RVSA 86 97.03/98.50 93.93/95.69 75.80 68.06 51.95 -
+LSKNet [44] LSKNet-S 14 97.05/98.22 94.27/95.83 - - 54.00 -
+OREOLE [18] ViT-G 914 96.71/   -    -   /   - 77.40 71.31 54.00 -
+SelectiveMAE* ViT-B 86 96.78/98.12 93.35/94.58 75.70 67.78 53.05 79.50
+SelectiveMAE* ViT-L 307 97.25/98.48 94.57/95.77 77.80 70.31 54.31 79.46
+SelectiveMAE ViT-B 86 96.64/98.30 92.53/94.50 75.10 67.69 52.68 79.44
+SelectiveMAE ViT-L 307 97.21/98.48 94.16/95.90 78.70 71.75 53.92 79.48
 4.2 Main Results
-4.2.1 Fine-tuning Performance on Downstream Tasks
-We further evaluated SelectiveMAE with the OpticalRS-4M dataset across three key downstream tasks: scene classification, object detection, and semantic segmentation.
-In addition to benchmarking against MAE-based approaches like SatMAE[30], ScaleMAE[33], SSL4EO[54], RVSA [18], RingMo[37], and GFM[25], we compared it with other leading visual foundation models.
-These encompass methods leveraging contrastive learning and enhanced backbones, such as GASSL [24], SeCo [22], TOV [28], and CACo [21], and advanced supervised learning-based approaches [56].
-Notably, none of these comparative methods were specifically tailored for accelerated pre-training, and although their speed performance wasn’t reported, we anticipate they may lag behind the computationally efficient baseline (MAE).
-More details can be found in the Appendix.
+We compare SelectiveMAE to state-of-the-art RSFMs.
+In addition to benchmarking against MIM-based approaches such as SatMAE [17] and ScaleMAE [56], we also compare with contrastive learning-based methods, including GASSL [1] and SeCo [49], as well as advanced supervised learning models like SatLas [5].
+Notably, none of these comparative methods were specifically designed for accelerated pre-training.
+Although their speed performance was not reported, we expect them to lag behind the baseline MAE.
 Scene Classification.
-We begin by evaluating the pre-trained model’s performance on the scene classification task, which offers insight into its overall representation capability without requiring additional decoders.
-We leverage two scene classification datasets: AID [85] and RESISC-45 [86].
-Training details, including the training and test split ratio, adhere to [18, 33], with further specifics available in our code.
-Evaluation is based on overall accuracy (OA).
-Table 6 presents the results, demonstrating SelectiveMAE’s competitive performance compared to other pre-training methods on both datasets.
+We first assess the pre-trained model’s performance on the scene classification task to evaluate its overall representational capability without additional decoders.
+We use two scene classification datasets: AID [86] and RESISC-45 [13], following training details and train-test splits as in [70, 56].
+Table 5 shows that SelectiveMAE performs competitively against other pre-training methods on both datasets.
 Notably, our method enjoys a significantly faster training speed, e.g., over 2$\times$ faster than the MAE baseline.
-Besides, it outperforms RVSA and others when scaled to ViT-L.
-This demonstrates that, with our SelectiveMAE training method on the OpticalRS-4M dataset, the model can learn strong feature representations and scale efficiently.
-Table 6: Results for scene classification, object detection, and semantic segmentation. “TR" represents the ratio of training data to the entire dataset.
-Model Publication Backbone Sence Object Semantic
-Classification Detection Segmentation
-AID RESISC-45 DIOR DIOR-R LoveDA SpaceNetv1
-(TR=20%/50%) (TR=10%/20%) mAP50 mAP50 mIoU mF1
-SeCo [22] ICCV’21 ResNet-50 93.47/95.99 89.64/92.91 - - 43.63 77.09
-GASSL [24] ICCV’21 ResNet-50 93.55/95.92 90.86/93.06 67.40 65.65 48.76 78.51
-TOV [28] JSTARS’23 ResNet-50 95.16/97.09 90.97/93.79 70.16 66.33 49.70 -
-CACo [21] CVPR’23 ResNet-50 90.88/95.05 88.28/91.94 66.91 64.10 48.89 77.94
-SatMAE [30] NIPS’22 ViT-L 95.02/96.94 91.72/94.10 70.89 65.66 - 78.07
-ScaleMAE [33] ICCV’23 ViT-L 96.44/97.58 92.63/95.04 73.81 66.47 - -
-SSL4EO [54] GRSM’23 ViT-S 91.06/94.74 87.60/91.27 64.82 61.23 - -
-RingMo [37] TGRS’22 Swin-B 96.90/98.34 94.25/95.67 75.90 - - -
-SatLas [56] ICCV’23 Swin-B 94.96/97.38 92.16/94.70 74.10 67.59 - -
-GFM [25] ICCV’23 Swin-B 95.47/97.09 92.73/94.64 72.84 67.67 - -
-RVSA [18] TGRS’23 ViT-B+RVSA 97.03/98.50 93.93/95.69 75.80 68.06 51.95 -
-SelectiveMAE - ViT-B 96.78/98.12 93.35/94.58 75.70 67.78 53.05 79.50
-SelectiveMAE - ViT-L 97.25/98.48 94.57/95.77 77.80 70.31 54.31 79.46
+Additionally, it surpasses LSKNet and others when scaled to ViT-L, indicating that, with SelectiveMAE on the OpticalRS-13M dataset, the model effectively learns strong feature representations and scales efficiently.
 Horizontal & Oriented Object Detection.
-We utilized the well-established DIOR dataset for horizontal object detection [87] and its improved variant DIOR-R for oriented object detection [88], both comprising RGB images.
+We utilized the well-established DIOR dataset for horizontal object detection [41] and its improved variant DIOR-R for oriented object detection [15].
 Following the methodologies of previous works [70, 63], we maintained consistent experimental setups, employing Faster-RCNN [57] and Oriented-RCNN [88] as detectors for each dataset.
-Detailed experimental configurations are available in the accompanying code.
 The results are summarized in Table 5.
 Our approach, utilizing a ViT-B backbone, demonstrates competitive or superior performance compared to other methods with a Swin-B backbone, such as RingMo [63].
-When using the larger ViT-L backbone, SelectiveMAE shows enhanced performance across both detection datasets, underscoring the excellent scalability of our method.
+When using the larger ViT-L backbone, SelectiveMAE shows enhanced performance across both detection datasets, even outperforming OREOLE, which has close to 1B parameters, underscoring the excellent scalability of our method.
 Semantic Segmentation.
 We further evaluate the performance of the pre-trained model on pixel-level perception tasks, particularly semantic segmentation, using two well-known remote sensing datasets: LoveDA [76] and SpaceNetv1 [68].
-Our implementation closely follows [18], utilizing UperNet [91] as the segmentation framework, with mean Intersection over Union (mIOU) as the evaluation metric.
-Table 6 demonstrates the clear superiority of SelectiveMAE over its competitors in pixel-level semantic segmentation.
-The selection and reconstruction strategies employed by SelectiveMAE focus on semantically rich patches, such as the boundaries between foreground objects and background stuff, leading to better representation learning for segmentation.
+Our implementation follows [70], utilizing UperNet [87] as the segmentation framework.
+Table 5 demonstrates the clear superiority of SelectiveMAE over its competitors in semantic segmentation tasks.
+SelectiveMAE focus on semantically rich patches, such as the boundaries between foreground objects and background stuff, leading to better representation learning for segmentation.
 5 Conclusion
 In this paper, we introduce a new pre-training pipeline for RS models, featuring the creation of a large-scale RS dataset and an efficient MIM approach.
-Built upon it, we first curated OpticalRS-4M, a large-scale optical remote sensing dataset for unsupervised learning.
-Unlike previous RS datasets, OpticalRS-4M offers a larger and more diverse image set with fine-grained details relevant to downstream tasks.
-Benchmarking representative MIM methods on OpticalRS-4M highlights its advantages in these tasks.
+Built upon it, we first curated OpticalRS-13M, a large-scale optical remote sensing dataset for unsupervised learning.
+Unlike previous RS datasets, OpticalRS-13M offers a larger and more diverse image set with fine-grained details relevant to downstream tasks.
+Benchmarking representative MIM methods on OpticalRS-13M highlights its advantages in these tasks.
 Then, we present SelectiveMAE to reduce the computational overhead of MIM training on large-scale RS datasets.
 This efficient MIM method dynamically encodes and reconstructs tokens based on their semantic richness.
-SelectiveMAE significantly accelerates training, demonstrating that only 40% RS image patches are needed to train a comparable MIM model.
-Extensive experiments demonstrate that OpticalRS-4M significantly improves classification, detection, and segmentation performance, while SelectiveMAE achieves over 2×2$\times$2 × speedup, highlighting the effectiveness and scalability of our pipeline in developing RS foundational models.
+SelectiveMAE significantly accelerates training, demonstrating that using only 40% RS image patches is sufficient for training a comparable MIM model.
+Extensive experiments show that OpticalRS-13M significantly contributes to improving classification, detection, and segmentation performance, while SelectiveMAE achieves over a 2×2$\times$2 × speedup along with GPU memory savings, highlighting the effectiveness and scalability of our pipeline in developing RS foundational models.

### v5

@@
-Scaling Efficient Masked Image Modeling on Large Remote Sensing Dataset
+Harnessing Massive Satellite Imagery with Efficient Masked Image Modeling
 Abstract
@@
 To enhance efficiency, we propose SelectiveMAE, a pre-training method that dynamically encodes and reconstructs semantically rich patch tokens, thereby reducing the inefficiencies of traditional MIM models caused by redundant background pixels in RS images.
-Extensive experiments demonstrate that OpticalRS-13M significantly improves classification, detection, and segmentation performance, while SelectiveMAE increases training efficiency over 2$\times$ times.
+Extensive experiments show that OpticalRS-13M significantly improves classification, detection, and segmentation performance, while SelectiveMAE increases training efficiency over 2$\times$ times.
 This highlights the effectiveness and scalability of our pipeline in developing RS foundational models.
+The dataset, source code, and trained models will be released at SelectiveMAE.
 1 Introduction
 Figure 1: Advantages of SelectiveMAE on finetuning performance and pre-training efficiency.
-The accuracy is evaluated on the DIOR detection dataset [41], while the training efficiency is measured on NVIDIA A100 GPUs, with different versions of ViT [19].
-Over the past decade, advancements in remote sensing (RS) technology and data acquisition have improved applications in ecosystem monitoring [54], natural disaster management [6], among others [32, 74].
-These applications rely on essential capabilities such as scene classification [30, 14], object detection [43], change detection [85], and semantic segmentation [93].
-However, each of these downstream tasks often requires substantial computational resources to learn task-specific feature representations and develop specialized models.
+The accuracy is evaluated on the DIOR detection dataset [1] with different versions of ViT [2].
+The training efficiency is measured by Data Throughput/Minute, i.e., the processed images per minute, on 8 $\times$×NVIDIA A100 GPUs.
+Over the past decade, advances in remote sensing (RS) technology and data acquisition have enhanced applications in ecosystem monitoring [3], natural disaster management [4], among others [5, 6].
+These rely on essential capabilities such as scene classification [7, 8], object detection [9], change detection [10], and semantic segmentation [11].
+However, each task typically demands substantial computational resources to train specialized models and learn task-specific feature representations.
 Figure 2: Advantages of OpticalRS-13M in data volumn and diversity.
-(a) The volumn of OpticalRS-13M surpasses at least 4$\times$ compared to existing RS datasets.
+(a) The volume of OpticalRS-13M surpasses at least 4$\times$ compared to existing RS datasets.
 (b) OpticalRS-13M includes significantly more sample categories than others, spanning both object and event types.
-(c) OpticalRS-13M provides more abundant feature representations compared to other large-scale RS datasets.
-Here, 1,000 samples are separately selected from different datasets, and features are extracted using a pre-trained ViT-B.
-Thanks to advances in self-supervised learning methods, such as Masked Image Modeling (MIM) techniques [28, 89], the pre-training of visual foundation models has seen remarkable improvements [90, 94, 91, 95, 73, 77].
-Consequently, remote sensing foundation models (RSFMs) have recently emerged, offering general feature representations and achieving outstanding performance across various remote sensing downstream tasks [70].
-However, two challenges persist in the development of RSFMs.
-(i) Compared to the ImageNet-21k [58] dataset, previous RS datasets [48, 50, 49, 62] contain fewer samples (approximately 1 million vs.
-14 million), which impedes the sufficient MIM training of large backbones.
-(ii) These datasets, which primarily focus on global scene semantics [48, 50, 49], lack the diversity and fine-grained information of RS scenarios encountered in downstream tasks.
-This limitation restricts the generalization of the learned representations.
+y-axis denotes the data volume in (a-b).
+(c) OpticalRS-13M provides more diverse feature patterns than other large-scale RS datasets.
+Here, 1,000 samples are separately selected from different datasets, and their features from a pre-trained ViT-B are visualized using t-SNE [12].
+Recent advances in self-supervised learning, particularly Masked Image Modeling (MIM) techniques [13, 14], have significantly improved the pre-training of visual foundation models [15, 16, 17, 18, 19, 20].
+This progress has led to the emergence of remote sensing foundation models (RSFMs), which provide general feature representations and achieve state-of-the-art performance across various remote sensing tasks [21].
+However, RSFMs still face two key challenges: (i) Compared to ImageNet-21k [22], existing RS datasets [23, 24, 25, 26] are significantly smaller (approximately 1 million vs.
+14 million samples), limiting the effective MIM training of large models; and (ii) these datasets primarily capture global scene semantics [23, 24, 25] but lack the diversity and fine-grained details essential for downstream tasks, hindering the generalization of learned representations.
 To address these challenges, taking an example of visible light RS imagery, we propose a new pipeline to collect, create, and efficiently process a large RS dataset.
@@
 Another approach is to reduce the number of visible patch tokens input into the vision encoder [46, 47, 48], speeding up feature extraction.
-However, conventional MIM approaches, such as the encoding-then-decoding procedure, overlook the unique characteristics of RS images, which typically feature sparse foreground pixels and dense backgrounds [70, 65].
+Conventional MIM approaches, following the encoding-then-decoding procedure, overlook the unique characteristics of RS images, which typically feature sparse foreground pixels and dense backgrounds [21, 31].
 This raises two key questions about how to efficiently conduct MIM training in the RS field: 1) Is it necessary to reconstruct all the redundant background patches during the MIM decoding process?
@@
 For the second question, the intuition is that the patch tokens used in the encoding-then-decoding procedure should effectively capture feature dependencies in RS images.
-Regarding the above issues, in the second part of the pipeline, we introduce a MIM method called SelectiveMAE for efficiently processing RS images, which dynamically encodes and reconstructs patch tokens based on their semantic richness.
+Regarding the above issues, in the second part of the pipeline, we introduce an MIM method called SelectiveMAE for efficiently processing RS images, which dynamically encodes and reconstructs patch tokens based on their semantic richness.
 Specifically, SelectiveMAE utilizes the Histogram of Oriented Gradients (HOG) algorithm to quantify the semantic richness of patches.
@@
 In summary, our main contributions are as follows:
-1) We introduce a new pipeline to collect, create, and efficiently process a large optical RS dataset for developing RS foundation models.
-2) Using this pipeline, we create the OpticalRS-13M dataset, which is a large-scale RS dataset comprising 13 million optical images with diversified coverage scenarios.
-3) We introduce SelectiveMAE, an efficient MIM method tailored for RS image pre-training.
+  1. 1.
+    We introduce a new pipeline to collect, create, and efficiently process a large optical RS dataset for developing RS foundation models.
+Using this pipeline, we create the OpticalRS-13M dataset, which is a large-scale RS dataset comprising 13 million optical images with diversified coverage scenarios.
+  2. 2.
+    For this pipeline, we introduce SelectiveMAE, an efficient MIM method tailored for RS image pre-training.
 It significantly accelerates convergence and enhances representation learning compared to the original MIM approach.
-Experiment results suggest the effectiveness and scalability of the proposed pipeline.
+  3. 3.
+    Experiment results demonstrate the effectiveness and scalability of the proposed pipeline.
 OpticalRS-13M significantly enhances the performance of RS foundation models in downstream tasks, while SelectiveMAE achieves over 2×2$\times$2 × speedup in pre-training compared to MAE [13].
@@
 However, these datasets primarily target scene classification and often overlook fine-grained target information, limiting their utility for various downstream tasks.
-To address this gap, with the proposed pipeline, we introduce the OpticalRS-13M dataset, which is larger and more diverse, enhancing performance across multiple downstream tasks.
+To address this gap, we introduce the OpticalRS-13M dataset, which is larger and more diverse.
+Using this dataset for pre-training, it significantly enhances performance across multiple downstream tasks.
 Masked Image Modeling.
@@
 Self-supervised learning methods have recently been employed to extract representations from unlabeled RS data.
-Due to the inefficiency of designing pretext tasks and gathering required data for contrastive self-supervised methods [1, 50, 49, 81], recent advancements have primarily centered around generative self-supervised methods, especially in MIM.
+Due to the inefficiency of designing pretext tasks and gathering required data for contrastive self-supervised methods [27, 24, 25, 79, 80], recent advancements have primarily centered around generative self-supervised methods, especially in MIM.
 Specifically, many studies aim to improve generative self-supervised algorithms by leveraging general image knowledge [28, 29, 31], scaling up parameter sizes [32, 43], integrating spatio-temporal information [33, 34, 36], encompassing geometric attributes [21, 81], handling multi-sensor data [82, 83, 84, 85, 86], and employing multi-scale concepts [37, 39, 40].
 However, these methods have not effectively addressed the substantial computational burden associated with self-supervised pre-training in RS.
-In the paper, we propose a new pipeline that can collect, create, and efficiently process large amounts of optical RS data for developing RS foundation models, enhancing the practicality of MIM pre-training on large-scale datasets.
+In the paper, we propose a new pipeline that can collect, create, and efficiently process large amounts of optical RS data for developing RSFMs, enhancing the practicality of MIM pre-training on large-scale datasets.
+Figure 3: The comparison between OpticalRS-13M and existing RS large-scale datasets in terms of data volume and image type.
 3 Method
@@
 Recent progress in self-supervised pre-training for RSFMs is limited by the relatively small scale and diversity of existing RS datasets compared to natural scene datasets.
-To overcome this, we curate a large-scale RS dataset, OpticalRS-13M, through extensive data collection and preprocessing.
-Data Collection We began by reviewing publicly available RS datasets from the past decade.
-Recognizing that not all data are suitable for self-supervised pre-training, we adopted specific criteria for data collection, as outlined in DiRS [48]: 1) Diversity: A dataset is diverse if its images capture various typical visual features of relevant scenes or targets and offer complementarity.
-Intra-class Diversity: High within-class diversity ensures comprehensive representation of real-world target distributions.
-We include data from different sources of the same scene or target, incorporating change detection data to capture object variations by geographic location and imaging time.
-Inter-class Diversity: To enable self-supervised pretraining to distinguish categories effectively, we included as many fine-grained categories as possible, especially those with high semantic overlap, helping the model learn robust feature representations.
-2) Richness: Beyond diversity, richness is essential.
-We ensure varied content features and large sample sizes by collecting images under multiple conditions (e.g., weather, seasons, lighting), introducing variability in translation, viewpoint, object posture, spatial resolution, background, and occlusion.
-Sampling across imaging conditions provides a comprehensive real-world representation, enhancing model’s representation and generalization capabilities.
-3) Scalability: Scalability is crucial given the evolving applications of remote sensing images.
-We continuously expand the dataset based on scene and target categories across classification, detection, and segmentation tasks.
-In summary, following the guidelines, we constructed a large-scale RS image dataset with diverse coverage scenarios.
-Data Preprocessing Following data collection, challenges such as inconsistent data sources, oversized images, and redundant pixels remained.
-To address these, we implemented a standardized, scalable data preprocessing workflow consisting of four steps:
-1) We focused exclusively on visible light images for this study, excluding multispectral and SAR data, though future updates will incorporate these modalities for multi-modal self-supervised pre-training.
-2) To manage the large image sizes, we randomly cropped high-resolution images into smaller slices.
-3) We filtered out images with fewer than 64 or more than 1024 pixels.
-4) The remaining images were combined and duplicates were removed using a two-phase approach: first, a coarse phase with perceptual hashing [55], and then a refined phase involving manual review.
-This ensured that only highly similar images were excluded.
-As a result of these steps, we created the OpticalRS-13M dataset, which contains approximately 13 million high-quality RS images, significantly larger than previous representative RS datasets, as illustrated in Fig.
-2 left.
-In summary, we followed the DiRS principles for data collection, applying exclusion, slicing, and deduplication to ensure high-quality.
-We’ve screened and processed most small and medium-scale datasets from the past decade in this field.
-Dataset Diversity To demonstrate the broad scope of scenes and target categories covered by our dataset, we conducted a statistical analysis accompanied by a detailed label explanation, as shown in Fig.
-2 right.
+To overcome this issue, we curate a large-scale RS dataset with diverse coverage scenarios, named OpticalRS-13M.
+Detailed data collection and preprocessing are presented in the supplementary material.
+Figure 4: Overview of SelectiveMAE. It inputs fewer visible patches and reconstructs only partial patches to accelerate training efficiency.
+Data Volume.
+OpticalRS-13M comprises 13 million high-resolution visible light RS images, establishing a significant scale advantage over existing datasets, as illustrated in Figure 3.
+OpticalRS-13M contains 13,203,698 images with a total of 2,630,362,174,503 pixels, averaging 199,214 pixels per image.
+We also offer a high-quality, lightweight version: OpticalRS-4M.
+OpticalRS-4M contains 3,920,829 images with a total of 688,459,799,359 pixels, averaging 175,590 pixels per image.
+While large-scale datasets such as SSL4EO-S12 [58] and SatLasPretrain [60] incorporate diverse data formats, e.g., including multispectral imagery, OpticalRS-13M is uniquely focused on visible light data.
+The dataset is carefully curated from multi-sensor acquisitions, leveraging imagery from high-resolution satellite platforms such as WorldView [87], QuickBird [88], GeoEye [89], and others.
+This specialization positions OpticalRS-13M as the most extensive visible light RS image repository to date, offering unparalleled utility for training vision foundation models for RS applications.
+Dataset Diversity.
+To demonstrate the broad scope of scenes and target categories covered by our dataset, we conducted a statistical analysis accompanied by a detailed label explanation, as shown in Fig.
+2 (b) and (c).
 The analysis reveals that OpticalRS-13M comprises 12 main categories, each containing numerous subclasses.
-Unlike existing self-supervised optical remote sensing datasets, OpticalRS-13M not only offers more comprehensive scene and target information but also introduces “Events” as a distinct high-level category.
+Unlike existing self-supervised optical RS datasets, OpticalRS-13M not only offers more comprehensive scene and target information but also introduces “Events” as a distinct high-level category.
 This category includes labels such as “Fire”, “Flood”, “Landslide”, “Post-Earthquake”, and so on.
@@
 3.2.1 Preliminaries of MAE
-Figure 3: Overview of SelectiveMAE. It inputs fewer visible patches and reconstructs only partial patches to accelerate training efficiency.
 1) Masking.
@@
 MAE predicts the pixel values for each masked patch, with each element in the decoder output representing a patch’s pixel value vector.
-The loss function computes the mean squared error (MSE) between the reconstructed targets and original patches.
+The loss function computes the mean squared error between the reconstructed targets and original patches.
 However, applying MAE for self-supervised pre-training results in considerable training costs in terms of time complexity, especially when working with large RS datasets.
@@
 How can we achieve acceleration at a high masking ratio (e.g., 85%) while ensuring MAE completes the pre-training task for RS images?
-Inspired by the curriculum learning  [78, 7, 26], which follows the principle of learning from easy to hard, we introduce the Progressive Semantic Token Selection (PSTS) module for patch selection, as depicted in Fig.
-3.
-In this module, we begin by selecting a limited number of patches and then select additional patches based on them in the training epoch, dynamically transitioning from easily learned, semantically similar patches to more challenging, complementary ones.
+Inspired by curriculum learning  [90, 91, 92], which follows the principle of learning from easy to hard, we introduce the Progressive Semantic Token Selection (PSTS) module for patch selection, as depicted in Fig.
+4.
+In this module, we begin by selecting a limited number of patches and then select additional patches based on the similarity measure in the training epoch, dynamically transitioning from easily learned, semantically similar patches to more challenging, complementary ones.
 For initialization, we employ a HOG selection strategy to choose the initial patch set from $S^{N}=\left\{x_{i}^{p}\right\}_{i=1}^{N}$, with a proportion $s\in[0,(1-m)/2]$.
+$m$ is the mask ratio.
 We define the initial set of selected token as:
@@
 In contrast, if a token’s feature distribution differs significantly from others, we consider these more challenging patches to learn.
+Figure 6: Illustration of the patch selection process in PSTS.
 Algorithm 1 Progressive Semantic Token Selection
@@
 We use Cosine Distance to measure the dissimilarity between the tokens in these two sets:
-$\small\mathcal{D}(S^{U},S^{I})=\mathbf{1}-\cos\left\langle S^{U},S^{I}\right\rangle=\mathbf{1}-\frac{\mathbb{S^{U}}\mathbb{S^{I}}^{\mathrm{T}}}{\left|\kern-0.96873pt\left|\mathbb{S^{U}}\right|\kern-0.96873pt\right|\cdot\left|\kern-0.96873pt\left|\mathbb{S^{I}}\right|\kern-0.96873pt\right|},$ (3)
+$\mathcal{D}(\mathbf{S}^{U},\mathbf{S}^{I})=1-\cos\left\langle\mathbf{S}^{U},\mathbf{S}^{I}\right\rangle=1-\frac{\mathbf{S}^{U}(\mathbf{S}^{I})^{\mathrm{T}}}{\lVert\mathbf{S}^{U}\rVert\cdot\lVert\mathbf{S}^{I}\rVert}$ (3)
 where $\mathbf{1}$ is an all-one matrix.
@@
 […]
-$\small{S^{K}}={S^{I}}\cup\mit{S}^{*},\quad{S^{U}}={S^{U}}\setminus\mit{S}^{*},$ (6)
+${S^{K}}={S^{I}}\cup\mit{S}^{*},\quad{S^{U}}={S^{U}}\setminus\mit{S}^{*},$ (6)
 where $\mit{S}^{*}$ represents the selected token from ${S^{U}}$.
@@
 6 further provides an example to illustrate which patches should be selected at different stages, helping to facilitate readers’ understanding.
-Figure 5: Illustration of the patch selection process in PSTS.
 4 Experiment
-Table 1: Performance comparison of using the MillionAID and OpticalRS-13M datasets separately with the same data throughput, evaluated by MAE [28] and ViT-B [19].
-Dataset Method Image Number Epoch Scene Object Semantic
-Classification Detection Segmentation
-AID [86] RESISC-45 [13] DIOR [41] DIOR-R [15] LoveDA [76] SpaceNetv1 [68]
-OA (TR=20% / 50%) OA (TR=10% / 20%) $\mathrm{mAP}_{50}$ $\mathrm{mAP}_{50}$ mIoU mF1
-MillionAID [48] MAE 1 million 800 94.92/97.38 89.20/93.60 71.80 62.33 51.24 79.24
-OpticalRS-13M MAE 1 million 800 96.26/97.98 91.53/93.88 75.90 66.46 53.10 79.35
-OpticalRS-13M MAE 2 million 400 96.64/98.10 91.80/94.31 73.90 65.95 52.86 79.37
-OpticalRS-13M MAE 3 million 267 96.67/98.18 92.24/94.41 75.40 67.07 52.39 79.37
-OpticalRS-13M MAE 4 million 200 96.10/98.03 92.38/94.30 74.70 66.26 52.75 79.23
-OpticalRS-13M MAE 8 million 100 96.58/98.26 91.83/93.99 75.20 66.99 52.56 79.43
-OpticalRS-13M MAE 13 million 67 96.28/98.06 91.41/93.60 74.50 66.36 52.63 79.39
-In this section, we first discuss the effectiveness of the dataset and acceleration strategy in the pipeline through ablation experiments.
-We also conducted further experiments to test the scalability of the pipeline in the ablation experiments.
-Additionally, we performed comprehensive comparative experiments to evaluate the pipeline’s performance on downstream tasks.
-All experiments are performed by fine-tuning pre-trained models on downstream tasks, including RS scene classification, object detection, and semantic segmentation.
+In this section, we performed comprehensive comparative experiments to evaluate the performance of the proposed method on various downstream tasks, including RS scene classification, object detection, and semantic segmentation.
+Then, we validated the effectiveness of the OpticalRS-13M dataset and the SelectiveMAE method through ablation experiments.
+We also conducted further experiments to test the scalability of the pipeline.
+All experiments are performed by fine-tuning.
 The pre-training and fine-tuning settings of our methods can be found in the supplementary material.
-Table 2: The performance comparison between SelectiveMAE and the baseline method MAE [28] with the ViT-B backbone.
-The training time is measured on NVIDIA A100 GPUs.
-To accelerate validation, we select 4 million images from OpticalRS-13M for pre-training.
-Method Image Number Epoch Total Training Scene Object Semantic
-Classification Detection Segmentation
-Time (h) AID [86] RESISC-45 [13] DIOR [41] DIOR-R [15] LoveDA [76] SpaceNetv1 [68]
-OA (TR=20% / 50%) OA (TR=10% / 20%) $\mathrm{mAP}_{50}$ $\mathrm{mAP}_{50}$ mIoU mF1
-MAE 4 million 800 177 96.88/98.22 92.44/94.43 75.40 67.35 52.80 79.41
-SelectiveMAE 4 million 800 86 (2.1×2.1$\times$2.1 ×) 96.78/98.12 93.35/94.58 75.70 67.78 53.05 79.50
-4.1 Ablation Study
-To validate the pipeline’s effectiveness, we conducted separate evaluations of both the OpticalRS-13M dataset generated by the pipeline and the accelerated training method, SelectiveMAE.
-Additionally, we assessed the pipeline’s scalability in collecting and processing datasets of varying sizes.
-More ablation experiment results related to model design are presented in the supplementary material.
-Dataset Diversity of OpticalRS-13M.
-In self-supervised pre-training of RSFMs, the primary goal is to enhance downstream task performance.
-Here, We define dataset diversity as the ability to improve performance across various types of tasks.
-While datasets like MillionAID [48] focus on classification, our OpticalRS-13M dataset is designed to extend the benefit to detection and segmentation tasks.
-We assess this by evaluating the effectiveness of models trained on different datasets with equivalent data throughput.
-As shown in Table 1, even with random sampling of 1 million images from OpticalRS-13M, it outperforms MillionAID on downstream tasks, particularly in object detection and segmentation.
-The results in Table 1 also indicate that the optimal outcomes at equivalent data throughput are achieved by different settings other than using all 13 million images.
-This suggests that training on a large number of samples with few epochs was insufficient for the MIM method to fully leverage the large dataset, leading to inferior generalization on downstream tasks.
-Experimental results in the supplementary material show consistent performance improvements across downstream tasks with more training epochs.
-Table 3: The efficiency comparison between SelectiveMAE and the baseline method MAE [28] across different backbones on the MillionAID dataset [48] with 1600 epochs.
-The memory is measured on a single NVIDIA A100 GPU with a batch size of 256.
-Model Backbone Params Total Training Time (h) GPU Memory (MB)
-MAE ViT-B 86M 101 17628
-SelectiveMAE ViT-B 86M 48 (2.1×2.1$\times$2.1 ×) 9570 (1.8×1.8$\times$1.8 ×)
-MAE ViT-L 307M 113 30530
-SelectiveMAE ViT-L 307M 50 (2.3×2.3$\times$2.3 ×) 18790 (1.6×1.6$\times$1.6 ×)
-Performance Advantage of SelectiveMAE.
-To further validate the accelerated pre-training strategy of our pipeline, we conducted extensive experiments on OpticalRS-13M using both MAE [28] and SelectiveMAE.
-The performance of ViT-B on various downstream tasks after pre-training is presented in Table 2.
-The results indicate that after pre-training on OpticalRS-13M, SelectiveMAE not only outperforms MAE but also achieves this at 2.1 times the speed.
-This makes SelectiveMAE particularly well-suited for large-scale datasets, offering significant time savings during training on OpticalRS-13M.
-Efficiency Advantage of SelectiveMAE.
-To further highlight the efficiency advantage of SelectiveMAE, we evaluate the training time and memory footprint during pretraining of different backbones on the MillionAID dataset, as shown in Table 3 (the corresponding accuracies are presented in Table 5).
-It can be seen that, since only 40% of the image patches (15% for the encoder and 25% for the decoder) are involved in the calculation, our methods present more than doubled training acceleration compared to the vanilla MAE baseline, where the acceleration advantage is more significant for larger models.
-Additionally, it is worth noting that SelectiveMAE also reduces memory footprint by using fewer patches.
-Scalability of the Pipeline To evaluate the scalability of our pipeline, we organized datasets containing 1 million and 4 million samples and conducted experiments on each.
-The results in Table 4 show that with 1 million samples, the proposed SelectiveMAE method reduced training time by 29 hours.
-When the dataset was expanded to 4 million samples, the time savings increased to 81 hours.
-These findings demonstrate that our pipeline scales effectively with larger datasets, offering significant pre-training acceleration.
-Table 4: The training cost of SelectiveMAE and the baseline method MAE [28] on OpticalRS-13M with different data volumes. All experiments are conducted on NVIDIA A100 GPUs.
-Method Images Number Epoch Total Training Time (h)
-MAE 1 million 800 54
-SelectiveMAE 1 million 800 25 (-29)
-MAE 4 million 800 177
-SelectiveMAE 4 million 800 86 (-81)
 Table 1: Performance comparison results of different models.
@@
 The first and second scores are marked by bold and blue, respectively.
-* means pre-training on 4 million images sampled from OpticalRS-13M.
-Model Backbone Params (M) Scene Object Semantic
-Classification Detection Segmentation
-AID [86] RESISC-45 [13] DIOR [41] DIOR-R [15] LoveDA [76] SpaceNetv1 [68]
-(TR=20%/50%) (TR=10%/20%) mAP50 mAP50 mIoU mF1
-SeCo [49] ResNet-50 26 93.47/95.99 89.64/92.91 - - 43.63 77.09
-GASSL [1] ResNet-50 26 93.55/95.92 90.86/93.06 67.40 65.65 48.76 78.51
-TOV [65] ResNet-50 26 95.16/97.09 90.97/93.79 70.16 66.33 49.70 -
-CACo [50] ResNet-50 26 90.88/95.05 88.28/91.94 66.91 64.10 48.89 77.94
-SatMAE [17] ViT-L 307 95.02/96.94 91.72/94.10 70.89 65.66 - 78.07
-ScaleMAE [56] ViT-L 307 96.44/97.58 92.63/95.04 73.81 66.47 - -
-SSL4EO [79] ViT-S 22 91.06/94.74 87.60/91.27 64.82 61.23 - -
-RingMo [63] Swin-B 88 96.90/98.34 94.25/95.67 75.90 - - -
-SatLas [5] Swin-B 88 94.96/97.38 92.16/94.70 74.10 67.59 - -
-GFM [51] Swin-B 88 95.47/97.09 92.73/94.64 72.84 67.67 - -
-RVSA [70] ViT-B+RVSA 86 97.03/98.50 93.93/95.69 75.80 68.06 51.95 -
-LSKNet [44] LSKNet-S 14 97.05/98.22 94.27/95.83 - - 54.00 -
-OREOLE [18] ViT-G 914 96.71/   -    -   /   - 77.40 71.31 54.00 -
-SelectiveMAE* ViT-B 86 96.78/98.12 93.35/94.58 75.70 67.78 53.05 79.50
-SelectiveMAE* ViT-L 307 97.25/98.48 94.57/95.77 77.80 70.31 54.31 79.46
-SelectiveMAE ViT-B 86 96.64/98.30 92.53/94.50 75.10 67.69 52.68 79.44
-SelectiveMAE ViT-L 307 97.21/98.48 94.16/95.90 78.70 71.75 53.92 79.48
+$\dagger$ means pre-training on 4 million images sampled from OpticalRS-13, with the epoch is set to 800.
+The overall accuracy is adoped as the metric for the scene classification task.
+Model Backbone Params (M) Data Scene Object Semantic
+Throughput Classification Detection Segmentation
+/Minute AID [93] RESISC-45 [94] DIOR [1] DIOR-R [95] LoveDA [55] SpaceNetv1 [56]
+TR=20%/50% TR=10%/20% mAP50 mAP50 mIoU mF1
+SeCo [25] ResNet-50 [96] 26 - 93.47/95.99 89.64/92.91 - - 43.63 77.09
+GASSL [27] ResNet-50 [96] 26 - 93.55/95.92 90.86/93.06 67.40 65.65 48.76 78.51
+TOV [31] ResNet-50 [96] 26 - 95.16/97.09 90.97/93.79 70.16 66.33 49.70 -
+CACo [24] ResNet-50 [96] 26 - 90.88/95.05 88.28/91.94 66.91 64.10 48.89 77.94
+SatMAE [33] ViT-L [2] 307 205k 95.02/96.94 91.72/94.10 70.89 65.66 - 78.07
+ScaleMAE [37] ViT-L [2] 307 206k 96.44/97.58 92.63/95.04 73.81 66.47 - -
+SSL4EO [58] ViT-S [2] 22 - 91.06/94.74 87.60/91.27 64.82 61.23 - -
+RingMo [41] Swin-B [97] 88 - 96.90/98.34 94.25/95.67 75.90 - - -
+SatLas [60] Swin-B [97] 88 243k 94.96/97.38 92.16/94.70 74.10 67.59 - -
+GFM [28] Swin-B [97] 88 - 95.47/97.09 92.73/94.64 72.84 67.67 - -
+RVSA [21] ViT-B+RVSA [21] 86 - 97.03/98.50 93.93/95.69 75.80 68.06 51.95 -
+OREOLE [43] ViT-G [43] 914 - 96.71/   -    -   /   - 77.40 71.31 54.00 -
+MAE [13$\dagger$ ViT-B [2] 86 264k 96.58/98.02 92.44/94.43 75.40 67.35 52.80 79.41
+SelectiveMAE $\dagger$ ViT-B [2] 86 556k 96.90/98.12 93.35/94.58 75.70 67.78 53.05 79.50
+[…]
+SelectiveMAE ViT-B [2] 86 556k 97.10/98.28 93.70/95.48 75.80 67.69 52.68 79.44
+SelectiveMAE ViT-L [2] 307 533k 97.49/98.52 94.73/96.36 78.70 71.75 53.92 79.48
 4.1 Main Results
@@
 In addition to benchmarking against MIM-based approaches such as SatMAE [33] and ScaleMAE [37], we also compare with contrastive learning-based methods, including GASSL [27] and SeCo [25], as well as advanced supervised learning models like SatLas [60].
-Notably, none of these comparative methods were specifically designed for accelerated pre-training.
+Notably, these comparative methods were specifically designed for increasing performance rather than accelerating pre-training.
+The Scene Classification and Detection comparison results are taken directly from SkySense [83], while those for Semantic Segmentation are exactly sourced from MTP [38].
 Although their speed performance was not reported, we expect them to lag behind the baseline MAE.
+The performance of ViT-B and ViT-L on various downstream tasks after pre-training is presented in Table 1.
+The results indicate that after pre-training on OpticalRS-13M, SelectiveMAE outperforms other RSFMs.
+Notably, it not only outperforms baseline (MAE) but also achieves this at 2.1 times the speed (556/264).
+This makes SelectiveMAE particularly well-suited for large-scale datasets, offering significant time savings during training on OpticalRS-13M.
 Scene Classification.
@@
 Table 1 shows that SelectiveMAE performs competitively against other pre-training methods on both datasets.
-Notably, our method enjoys a significantly faster training speed, e.g., over 2$\times$ faster than the MAE baseline.
-Additionally, it surpasses LSKNet and others when scaled to ViT-L, indicating that, with SelectiveMAE on the OpticalRS-13M dataset, the model effectively learns strong feature representations and scales efficiently.
+Additionally, when scaled to ViT-L, our model outperforms OREOLE [43] and other competitors, indicating the efficacy of SelectiveMAE pre-training on the OpticalRS-13M dataset, which enables strong feature representation learning while maintaining efficient scalability with increasing model size.
 Horizontal & Oriented Object Detection.
@@
 Our approach, utilizing a ViT-B backbone, demonstrates competitive or superior performance compared to other methods with a Swin-B backbone, such as RingMo [41].
-When using the larger ViT-L backbone, SelectiveMAE shows enhanced performance across both detection datasets, even outperforming OREOLE, which has close to 1B parameters, underscoring the excellent scalability of our method.
+When using the larger ViT-L backbone, SelectiveMAE shows enhanced performance across both detection datasets, even outperforming OREOLE [43], which has close to 1B parameters, underscoring the excellent scalability of our method.
 Semantic Segmentation.
-We further evaluate the performance of the pre-trained model on pixel-level perception tasks, particularly semantic segmentation, using two well-known remote sensing datasets: LoveDA [76] and SpaceNetv1 [68].
+We further evaluate the performance of the pre-trained model on pixel-level perception tasks, particularly semantic segmentation, using two well-known RS datasets: LoveDA [55] and SpaceNetv1 [56].
 Our implementation follows [21], utilizing UperNet [100] as the segmentation framework.
@@
 SelectiveMAE focus on semantically rich patches, such as the boundaries between foreground objects and background stuff, leading to better representation learning for segmentation.
+4.2 Ablation Study
+To validate the efficacy of SelectiveMAE, we conducted ablation experiments about HOG Selecting, Similarity Measure and Selection Pattern.
+To further verify the pipeline’s effectiveness, we conducted separate evaluations of both the OpticalRS-13M dataset generated by the pipeline and the accelerated performance of SelectiveMAE.
+Additionally, we compared the SelectiveMAE with the CrossMAE, a MAE-based method for natural images.
+More ablation experiment results are presented in the supplementary material.
+Alternatives of HOG Selecting.
+To verify the efficacy of HOG in token selecting, we conduct two ablation experiments using MillionAID dataset [23] to replace HOG: (1) we used a pretrained Swin-B [97] model and applied k-nearest neighbor clustering to the extracted features.
+(2) We employed a lightweight feature extractor network based on Adamae [101].
+Results in the Table 2 demonstrate HOG’s pre-training speed advantage.
+While using Swin-B slightly outperforms HOG, it is highly inefficient, and pre-training a foundation model like Swin requires massive data and computational resources.
+For RS images with abundant low-level features, HOG effectively selects tokens while ensuring substantially higher training speeds.
+Table 2: Ablation experiment results of HOG selection through pre-training on the MillionAID dataset for 800 epochs.
+Method Params Data Throughput AID RESISC-45
+/ Minute TR=20%/50% TR=10%/20%
+Adamae [101] 2.36M 498k 88.78/91.25 85.72/87.44
+Swin-B 88M 356k 93.21/96.48 89.94/93.72
+HOG - 556k 93.17/96.12 89.21/92.31
+Ablation of Similarity Measure and Selection Pattern.
+To comprehensively assess the PSTS, we tested different Similarity Measures and the Selection Pattern.
+The results in the Table 3 reveal that different similarity measures had minimal impact on performance.
+However, changing the selection pattern led to a performance drop, confirming the effectiveness of the PSTS.
+Table 3: Ablation experiment results of different similarity measures and selection patterns in PSTS through pre-training on the MillionAID dataset for 800 epochs.
+Selection Pattern Similarity Measure AID RESISC-45
+TR=20%/50% TR=10%/20%
+near-far-random Euclidean distance 93.02/95.96 88.97/92.07
+near-far-random Manhattan distance 92.92/95.89 89.17/91.92
+far-near-random Cosine distance 90.12/92.78 85.81/88.80
+near-far-random Cosine distance 93.17/96.12 89.21/92.31
+Table 4: Performance comparison of using different pre-training settings on datasets, methods and epochs, where the ViT-B [2] is adopted as the network.
+Dataset Method Image Number Epoch AID [93] RESISC-45 [94]
+TR=20%/50% TR=10%/20%
+Different Datasets
+MillionAID [23] MAE 1 million 800 94.92/97.38 89.20/93.60
+OpticalRS-13M MAE 1 million 800 96.26/97.98 91.53/93.88
+Equivalent Data Throughput
+OpticalRS-13M MAE 2 million 400 96.64/98.10 91.80/94.31
+OpticalRS-13M MAE 3 million 267 96.67/98.18 92.24/94.41
+OpticalRS-13M MAE 4 million 200 96.10/98.03 92.38/94.30
+OpticalRS-13M MAE 8 million 100 96.58/98.26 91.83/93.99
+OpticalRS-13M MAE 13 million 67 96.28/98.06 91.41/93.60
+OpticalRS-13M SelectiveMAE 1 million 800 96.29/97.78 91.41/93.48
+OpticalRS-13M SelectiveMAE 4 million 200 95.96/98.06 92.06/94.05
+OpticalRS-13M SelectiveMAE 13 million 67 96.31/97.95 91.88/93.76
+More Epoch
+OpticalRS-13M MAE 13 million 67 96.28/98.06 91.41/93.60
+OpticalRS-13M MAE 13 million 100 96.34/98.13 91.79/93.85
+OpticalRS-13M MAE 13 million 200 96.51/98.22 92.19/94.26
+OpticalRS-13M MAE 13 million 800 97.10/98.28 93.70/95.48
+Efficacy of OpticalRS-13M.
+As illustrated in Table 4, when pre-training with a randomly sampled subset of 1 million images from OpticalRS-13M, the model achieves superior performance compared to MillionAID [23], underscoring the efficacy of OpticalRS-13M for representation learning.
+To further assess dataset diversity, we conducted experiments with equivalent data throughput during pre-training.
+The results in Table 4 reveal that optimal performance is attained by varying training configurations rather than utilizing the entire 13 million image corpus, regardless of the pre-training methodology.
+In our opinion, this observation highlights the inherent diversity of OpticalRS-13M, as training with fewer epochs proved insufficient for the MIM method to fully exploit the dataset’s potential, leading to model underfitting.
+To validate this hypothesis, we extended the training schedule, and the experimental results presented in the last part of Table 4 confirm our claim.
+Table 5: The efficiency comparison between SelectiveMAE and the baseline method MAE [13] across different backbones on the part of OpticalRS-13M dataset [23] with 800 epochs.
+The memory is measured on a single NVIDIA A100 GPU with a batch size of 256.
+Model Backbone Data Volume Training Time (h) GPU Memory (MB)
+MAE ViT-B 1 million 51 17628
+SelectiveMAE ViT-B 1 million 24 (2.1×2.1$\times$2.1 ×, -27) 9570 (1.8×1.8$\times$1.8 ×)
+MAE ViT-L 1 million 57 30530
+SelectiveMAE ViT-L 1 million 25 (2.3×2.3$\times$2.3 ×, -32) 18790 (1.6×1.6$\times$1.6 ×)
+MAE ViT-B 4 million 177 17628
+SelectiveMAE ViT-B 4 million 86 (2.1×2.1$\times$2.1 ×, -81) 9570 (1.8×1.8$\times$1.8 ×)
+Efficiency Advantage of SelectiveMAE.
+To further highlight the efficiency advantage of SelectiveMAE, we evaluate the training time and memory footprint during pre-training of different backbones on the part of OpticalRS-13M dataset, as shown in Table 5 (the corresponding accuracies are presented in Table 1).
+It can be seen that, since only 40% of the image patches (15% for the encoder and 25% for the decoder) are involved in the calculation, our methods present more than doubled training acceleration compared to the vanilla MAE baseline, where the acceleration advantage is more significant for larger models.
+It is worth noting that SelectiveMAE also reduces memory footprint by using fewer patches.
+To further evaluate the scalability of our pipeline, we organized datasets containing 4 million samples and conducted experiments.
+The results in Table 5 show that, when the dataset was expanded to 4 million samples, the time savings increased from 27 (51-24) to 81 hours.
+These findings demonstrate that our pipeline scales effectively with larger datasets, offering significant pre-training acceleration.
+Comparison with CrossMAE [49].
+CrossMAE, an MAE-based self-supervised pretraining method, is originally designed for natural images.
+Since SelectiveMAE adopts the same decoder as CrossMAE [49] during pre-training, and incorporates a HOG-based strategy tailored to RS images inspired by its partial reconstruction concept, as detailed in Section 3.2.2.
+In this section, we have also complemented related comparison experiments.
+Nevertheless, due to CrossMAE has not been pretrained in the RS field, we conducted a fair evaluation through pretraining it on the OpticalRS-13M dataset using the official code.
+Table 6 shows that SelectiveMAE outperforms CrossMAE in both accuracy and efficiency.
+By integrating RS-specific enhancements, PSTS and partial reconstruction, SelectiveMAE proves to be a highly effective self-supervised learning approach for RS.
+Table 6: Performance comparison between our method and CrossMAE, where 400 million images of OpticalRS-13M are randomly sampled for pre-training 800 epochs.
+Method Data Throughput / Minute AID RESISC-45
+TR=20%/50% TR=10%/20%
+CrossMAE(ViT-B) 475k 96.12/97.18 92.08/94.12
+SelectiveMAE(ViT-B) 556k 96.78/98.12 93.35/94.58
 5 Conclusion
 In this paper, we introduce a new pre-training pipeline for RS models, featuring the creation of a large-scale RS dataset and an efficient MIM approach.
-Built upon it, we first curated OpticalRS-13M, a large-scale optical remote sensing dataset for unsupervised learning.
+We first curated OpticalRS-13M, a large-scale optical remote sensing dataset for unsupervised learning.
 Unlike previous RS datasets, OpticalRS-13M offers a larger and more diverse image set with fine-grained details relevant to downstream tasks.